### v2

# A Generalized Approach to Online Convex Optimization

Abstract
In this paper, we analyze the problem of online convex optimization in different settings. We show that any algorithm for online linear optimization with fully adaptive adversaries is an algorithm for online convex optimization. We also show that any such algorithm that requires full-information feedback may be transformed to an algorithm with semi-bandit feedback with comparable regret bound. We further show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. We use this to describe general meta-algorithms to convert first order algorithms to zeroth order algorithms with comparable regret bounds. Our framework allows us to analyze online optimization in various settings, such full-information feedback, bandit feedback, stochastic regret, adversarial regret and various forms of non-stationary regret, and recovers several results in the literature with a simplified proof technique.
1 Introduction
1.1 Overview
Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information Shalev-Shwartz, (2012); Hazan et al., (2016). This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$. The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, the information exchange through a query oracle.
In this paper, we present a comprehensive approach to solving online convex optimization problems, addressing scenarios where the adversary can choose a function from the class of $\mu$–strongly convex functions (or convex functions for $\mu=0$) at each time step. Our approach encompasses different feedback types, including bandit, semi-bandit and full-information feedback in each iteration. It also encompasses different types of regret, such as adversarial regret, stochastic regret, and various forms of non-stationary regret. While the problem has been studied in many special cases, this comprehensive approach sheds light on relations between different cases and the powers and limitations of some of the approaches used in different settings.
One of our key contribution lies in establishing that any semi-bandit feedback online linear (or quadratic) optimization algorithm for fully adaptive adversary is also an online convex (or strongly convex) optimization algorithm. While the above result is for semi-bandit feedback, we then show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough and more information is not needed. Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. Finally, we introduce a meta algorithm, based on the variant of classical result of Flaxman et al., (2005), that converts algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit feedback.
Figure 1: Summary of the main results
Each arrow describes a procedure or meta-algorithm to convert an algorithm from one setting into another. The first line describes the function class - linear/convex. The second lines specifies the type of the adversary - fully adaptive or oblivious. The last line starts with D or S, to denote deterministic or stochastic oracles. After that, SB is used to denotes semi-bandit, B to denote bandit and FI to denote full information feedback. Finally, the last D or S is used to denotes deterministic or stochastic algorithm.
1.2 Related works
The term “online convex optimization” was first defined in Zinkevich, (2003), where Online Gradient Descent (OGD) was proposed, while this setting was first introduced in Gordon, (1999). The Follow-The-Perturbed-Leader (FTPL) for online linear optimization was introduced by Kalai and Vempala, (2005), inspired by Follow-The-Leader algorithm originally suggested by Hannan, (1957). Follow-The-Regularized-Leader (FTRL) in the context of online convex optimization was coined in Shalev-Shwartz and Singer, (2007); Shalev-Shwartz, (2007). The equivalence of FTRL and Online Mirror Descent (OMD) was observed by Hazan and Kale, (2010). Non-stationary measures of regret have been studied in  Zinkevich, (2003); Hazan and Seshadhri, (2009); Besbes et al., (2015); Daniely et al., (2015); Zhang et al., 2018b ; Zhang et al., 2018a ; Zhao et al., (2020, 2021); Lu et al., (2023); Wang et al., (2024); Garber and Kretzu, (2022) and many other papers. We refer the reader to surveys Shalev-Shwartz, (2012) and Hazan et al., (2016) for more details.
1.3 Contributions
The key results of this paper are encapsulated in Figure 1, where each arrow represents a procedure or meta-algorithm that transforms an algorithm from one setting to another. The primary contribution is two-fold:
  • we provide a framework that simplifies analysis of meta-algorithms in online learning; and
  • using this framework, we formulate several meta-algorithms that facilitate the conversion of algorithms across various settings.
This framework not only simplifies proof analysis for numerous existing problems in the field (see Remark 2 and Corollaries 34 and 5) but also gives a tool for easier extension of multiple results in the literature. In the following, we go over some of the main meta-algorithms that we discuss in this paper.
The first key result in this paper is Theorem 2. This result shows that any algorithm for online linear optimization with stochastic subgradient feedback results in a comparable regret with if the function class is replaced by a class of convex functions. This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback. We note that this result recovers the result for online convex game in Abernethy et al., (2008), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
While the above result is the key for going from online linear to online convex optimization, the main issue for its applicability is that it requires semi-bandit feedback. In the remaining results, we provide meta-algorithms to increase the applicability of this result.
The second key result, given in Theorem 3, is that for any full-information algorithm for online convex optimization with fully adaptive adversary, there is a corresponding algorithm with regret bound of the same order which only requires semi-bandit feedback. This allows us to get results for semi-bandit feedback from that of full information. This combined with Theorem 2 allows us to go from any full-information algorithm for online linear optimization with fully adaptive adversary to an algorithm for online convex optimization with semi-bandit feedback and fully adaptive adversary. This combination has been marked as Theorem 3 in Figure 1.
We note that the above results explained in Theorems 14 assume that the feedback oracle is deterministic. In the next key result, given in Theorem 5, we relax this assumption. We show that any online optimization algorithm for fully adaptive adversaries using deterministic semi-bandit feedback obtains similar regret bounds when facing an oblivious adversary using only stochastic semi-bandit feedback. Thus, this result allows us to get the results for online convex optimization with stochastic semi-bandit feedback from the results with deterministic semi-bandit feedback, as can be seen in Fig. 1. This result gives Theorem 6.5 in Hazan et al., (2016) as a special case. Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
Finally, we note that the above discussion was mostly focused on full information and semi-bandit feedback settings. However, in many practical setups, zeroth order or bandit feedback is important. In the final key result, given in Theorems 6-7, we use the smoothing trick Flaxman et al., (2005); Pedramfar et al., (2023), in order to estimate the gradient from the value oracle. Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively. These results apply to any optimization problem, not necessarily convex. This approach recovers the results in Flaxman et al., (2005) with the base algorithm of OGD, while extends to say that the approach can be applied directly on any base algorithm. In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback Zhao et al., (2021); Garber and Kretzu, (2022) with a simplified proof.
Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings. For instance, we can use the results for a deterministic algorithm for online linear optimization with oblivious adversaries and deterministic semi-bandit feedback to that for a stochastic algorithm for online convex optimization with oblivious adversaries and stochastic bandit feedback using the different arrows as in Fig. 1. The key technical novelty is the recognition that many results, previously understood only in very limited contexts, actually apply more broadly. This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results. For instance, results in Abernethy et al., (2008); Hazan et al., (2016); Flaxman et al., (2005); Zhao et al., (2021) as mentioned above can be demonstrated as corollaries of these more generalized results.
2 Background and Notations
For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{K}}$ and $\alpha\in{\mathbb{R}}$. The relative interior of ${\mathcal{D}}$ is defined as
$\operatorname{relint}({\mathcal{D}}):=\{{\mathbf{x}}\in{\mathcal{D}}\mid\exists r>0,{\mathbb{B}}_{r}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{D}})\subseteq{\mathcal{D}}\}.$
A function class is simply a set of real-valued functions. Given a set ${\mathcal{D}}$, a function class over ${\mathcal{D}}$ is simply a subset of all real-valued functions on ${\mathcal{D}}$. Given a real number $\mu\geq 0$ and a set ${\mathcal{D}}$, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{D}}$. A set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is called a convex set if for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is called convex if, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $f(\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}})\leq\alpha f({\mathbf{x}})+(1-\alpha)f({\mathbf{y}})$. All convex functions are continuous on any point in the relative interior of their domains. In this work, we will only focus on continuous functions. If ${\mathbf{x}}\in\operatorname{relint}({\mathcal{K}})$ and $f$ is convex and is differentiable at ${\mathbf{x}}$, then we have $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle\nabla f({\mathbf{x}}),{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. A vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is called a subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex if and only if it has a subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$. More generally, given $\mu\geq 0$, we say a vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is a $\mu$-subgradient of $f$ at ${\mathbf{x}}$ if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, for all ${\mathbf{y}}\in{\mathcal{K}}$. Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is $\mu$-strongly convex if and only if it has a $\mu$-subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$.
3 Problem Setup
Online optimization problems can be formalized as a repeated game between an agent and an adversary. The game lasts for $T$ rounds on a convex domain ${\mathcal{K}}$ where $T$ and ${\mathcal{K}}$ are known to both players. In $t$-th round, the agent chooses an action ${\mathbf{x}}_{t}$ from an action set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$, then the adversary chooses a loss function $f_{t}\in{\mathbf{F}}$ and a query oracle for the function $f_{t}$. Then, for $1\leq i\leq k_{t}$, the agent chooses a points ${\mathbf{y}}_{t,i}$ and receives the output of the query oracle.
To be more precise, an agent consists of a tuple $(\Omega^{\mathcal{A}},{\mathcal{A}}^{\text{action}},{\mathcal{A}}^{\text{query}})$, where $\Omega^{\mathcal{A}}$ is a probability space that captures all the randomness of ${\mathcal{A}}$. We assume that, before the first action, the agent samples $\omega\in\Omega$. The next element in the tuple, ${\mathcal{A}}^{\text{action}}=({\mathcal{A}}^{\text{action}}_{1},\cdots,{\mathcal{A}}^{\text{action}}_{T})$ is a sequence of functions such that ${\mathcal{A}}_{t}$ that maps the history $\Omega^{\mathcal{A}}\times{\mathcal{K}}^{t-1}\times\prod_{s=1}^{t-1}({\mathcal{K}}\times{\mathcal{O}})^{k_{s}}$ to ${\mathbf{x}}_{t}\in{\mathcal{K}}$ where we use ${\mathcal{O}}$ to denote range of the query oracle. The last element in the tuple, ${\mathcal{A}}^{\text{query}}$, is the query policy. For each $1\leq t\leq T$ and $1\leq i\leq k_{t}$, ${\mathcal{A}}^{\text{query}}_{t,i}:\Omega^{\mathcal{A}}\times{\mathcal{K}}^{t}\times\prod_{s=1}^{t-1}({\mathcal{K}}\times{\mathcal{O}})^{k_{s}}\times({\mathcal{K}}\times{\mathcal{O}})^{i-1}$ is a function that, given previous actions and observations, either selects a point ${\mathbf{y}}_{t}^{i}\in{\mathcal{K}}$, i.e., query, or signals that the query policy at this time-step is terminated. We may drop $\omega$ as one of the inputs of the above functions when there is no ambiguity. We say the agent query function is trivial if $k_{t}=1$ and ${\mathbf{y}}_{t,1}={\mathbf{x}}_{t}$ for all $1\leq t\leq T$. In this case, we simplify the notation and use the notation ${\mathcal{A}}={\mathcal{A}}^{\text{action}}=({\mathcal{A}}_{1},\cdots,{\mathcal{A}}_{T})$ to denote the agent action functions and assume that the domain of ${\mathcal{A}}_{t}$ is $\Omega^{\mathcal{A}}\times({\mathcal{K}}\times{\mathcal{O}})^{t-1}$.
A query oracle is a function that provides the observation to the agent. Formally, a query oracle for a function $f$ is a map ${\mathcal{Q}}$ defined on ${\mathcal{K}}$ such that for each ${\mathbf{x}}\in{\mathcal{K}}$, the ${\mathcal{Q}}({\mathbf{x}})$ is a random variable taking value in the observation space ${\mathcal{O}}$. The query oracle is called a stochastic value oracle or stochastic zeroth order oracle if ${\mathcal{O}}={\mathbb{R}}$ and $f({\mathbf{x}})={\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$. Similarly, it is called a stochastic (sub)gradient oracle or stochastic first order oracle if ${\mathcal{O}}={\mathbb{R}}^{d}$ and ${\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$ is a (sub)gradient of $f$ at ${\mathbf{x}}$. In all cases, if the random variable takes a single value with probability one, we refer to it as a deterministic oracle. Note that, given a function, there is at most a single deterministic gradient oracle, but there may be many deterministic subgradient oracles. We will use $\nabla$ to denote the deterministic gradient oracle. We say an oracle is bounded by $B$ if its output is always within the Euclidean ball of radius $B$ centered at the origin. We say the agent takes semi-bandit feedback if the oracle is first-order and the agent query function is trivial. Similarly, it takes bandit feedback if the oracle is zeroth-order and the agent query function is trivial [^1] This is a slight generalization of the common use of the term bandit feedback. Usually, bandit feedback refers to the case where the oracle is a deterministic zeroth-order oracle and the agent query function is trivial.. If the agent query function is non-trivial, then we say the agent requires full-information feedback.
An adversary $\operatorname{Adv}$ is a set such that each element ${\mathcal{B}}\in\operatorname{Adv}$, referred to as a realized adversary, is a sequence $({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})$ of functions where each ${\mathcal{B}}_{t}$ maps a tuple $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})\in{\mathcal{K}}^{t}$ to a tuple $(f_{t},{\mathcal{Q}}_{t})$ where $f_{t}\in{\mathbf{F}}$ and ${\mathcal{Q}}_{t}$ is a query oracle for $f_{t}$  [^2] Note that we do not assign a probability to each realized adversary since the notion of regret simply computes the supremum over all realizations.. We say an adversary $\operatorname{Adv}$ is oblivious if for any realization ${\mathcal{B}}=({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})$, all functions ${\mathcal{B}}_{t}$ are constant, i.e., they are independent of $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})$. In this case, a realized adversary may be simply represented by a sequence of functions $(f_{1},\cdots,f_{T})\in{\mathbf{F}}^{T}$ and a sequence of query oracles $({\mathcal{Q}}_{1},\cdots,{\mathcal{Q}}_{T})$ for these functions. We say an adversary is a weakly adaptive adversary if each function ${\mathcal{B}}_{t}$ described above does not depend on ${\mathbf{x}}_{t}$ and therefore may be represented as a map defined on ${\mathcal{K}}^{t-1}$. In this work we also consider adversaries that are fully adaptive, i.e., adversaries with no restriction. Clearly any oblivious adversary is a weakly adaptive adversary and any weakly adaptive adversary is a fully adaptive adversary. Given a function class ${\mathbf{F}}$ and $i\in\{0,1\}$, we use $\operatorname{Adv}^{\text{f}}_{i}({\mathbf{F}})$ to denote the set of all possible realized adversaries with deterministic $i$-th order oracles. If the oracle is instead stochastic and bounded by $B$, we use $\operatorname{Adv}^{\text{f}}_{i}({\mathbf{F}},B)$ to denote such an adversary. Finally, we use $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}})$ and $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}},B)$ to denote all oblivious realized adversaries with $i$-th order deterministic and stochastic oracles, respectively.
In order to handle different notions of regret with the same approach, for an agent ${\mathcal{A}}$, adversary $\operatorname{Adv}$, compact set ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ and $1\leq a\leq b\leq T$, we define regret as
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ $:=\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\min_{{\mathbf{u}}=({\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{T})\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right],$
where the expectation in the definition of the regret is over the randomness of the algorithm and the query oracles. We use the notation ${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]:={\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ when $\operatorname{Adv}=\{{\mathcal{B}}\}$ is a singleton.
Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$. When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret Zinkevich, (2003); Zhang et al., 2018a . If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret Besbes et al., (2015). Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ Hazan and Seshadhri, (2009). We drop $a$, $b$ and ${\mathcal{U}}$ when the statement is independent of their value or their value is clear from the context.
Remark 1.
As a special case of static adversarial regret, if $\operatorname{Adv}$ is oblivious and every ${\mathcal{B}}\in\operatorname{Adv}$ corresponds to $f_{1}=f_{2}=\cdots=f_{T}=f$ and ${\mathcal{Q}}_{1}={\mathcal{Q}}_{2}\cdots={\mathcal{Q}}_{T}={\mathcal{Q}}$ for some $f\in{\mathbf{F}}$ and some stochastic oracle ${\mathcal{Q}}$ for $f$, then the regret is referred to as the stochastic regret.
Another metric for comparing algorithms is high probability regret bounds which we define for stochastic algorithms. Given $\omega\in\Omega$, we may consider ${\mathcal{A}}_{\omega}$ as a deterministic algorithm. We say $h:[0,1]\to[0,\infty)$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$ if for each ${\mathcal{B}}\in\operatorname{Adv}$ and $\delta\in[0,1]$, we have
(${\mathbb{P}}\left(\{\omega\in\Omega\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
Note that the infimum of any family of high probability regret bounds is a high probability regret bound. Hence we use ¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}}:[0,1]\to[0,\infty)$ to denote the smallest high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$. In other words, ¯$\overline{{\mathcal{R}}}_{{\mathcal{B}}}^{{\mathcal{A}}}$ is the quantile function for the random variable $\omega\mapsto{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}$.
4 Re-statement of Previous Result: Oblivious to fully adaptive adversary
The following theorem states a result that is well-known in the literature. For example, Hazan et al., (2016) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary. Here we explicitly mention it as a theorem for completion.
Theorem 1.
Let $i\in\{0,1\}$ and assume ${\mathcal{A}}$ is a deterministic online algorithm designed for $i$-th order feedback and ${\mathbf{F}}$ is a function class. Then we have
${\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}^{{\mathcal{A}}}.$
See Appendix A for proof.
Corollary 1.
The OGD, FTRL, and OMD algorithms are deterministic and therefore have same regret bound in fully-adaptive setting as the oblivious setting.
5 Linear to convex with fully adaptive adversary
Here we show that any semi-bandit feedback online linear optimization algorithm for fully adaptive adversary is also an online convex optimization algorithm. We start with a definition.
Definition 1.
Let ${\mathbf{F}}$ be a function class over ${\mathcal{K}}$ and let $\mu\geq 0$. We define ${\mathbf{F}}_{\mu}$, namely $\mu$-quadratization of ${\mathbf{F}}$, to be class of functions $q:{\mathcal{K}}\to{\mathbb{R}}$ such that there exists $f\in{\mathbf{F}}$, ${\mathbf{x}}\in{\mathcal{K}}$, and ${\mathbf{o}}\in\nabla^{*}f({\mathbf{x}})$ such that
$q({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}\in{\mathbf{Q}}_{\mu}.$
When $\mu=0$, we may also refer to ${\mathbf{F}}_{\mu}$ as the linearization of ${\mathbf{F}}$. We say ${\mathbf{F}}$ is closed under $\mu$-quadratization if ${\mathbf{F}}\supseteq{\mathbf{F}}_{\mu}$. Similarly, for $B>0$, we define ${\mathbf{F}}_{\mu}[B]$ to be the the class of functions $q$ defined above where instead we have ${\mathbf{o}}\in{\mathbb{B}}_{{\mathbf{0}}}(B)$.
Theorem 2.
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ be a convex set, let $\mu\geq 0$ and let ${\mathcal{A}}$ be algorithm for online optimization with semi-bandit feedback. If ${\mathbf{F}}$ be a $\mu$-strongly convex function class over ${\mathcal{K}}$, then we have
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
Moreover, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then we have equality.
See Appendix B for proof.
Corollary 2 (Theorem 14 in Abernethy et al., (2008)).
The min-max regret bound of the problem of online linear optimization with deterministic feedback and online convex optimization with deterministic feedback are equal.
To see why this is true, note that any algorithm has the same performance in both settings. So it follows that an optimal algorithm for one setting is optimal for the other and therefore the min-max regret bounds are equal.
6 Full information feedback to semi-bandit feedback
Any algorithm designed for semi-bandit setting may be trivially applied in the first-order full-information feedback setting. In the following theorem, we show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough. Specifically, we show that an algorithm that requires full information feedback could be converted to an algorithm that only requires semi-bandit feedback with the same regret bounds. The meta-algorithm that does this conversion is described in Algorithm 1. Next we show that $\operatorname{FTS}({\mathcal{A}})$ always performs at least as good as ${\mathcal{A}}$ when the oracle is deterministic.
Input :  horizon $T$, algorithm ${\mathcal{A}}=({\mathcal{A}}^{\text{action}},{\mathcal{A}}^{\text{query}})$, strong convexity coefficient $\mu\geq 0$
for $t=1,2,\dots,T$ do
       Play the action ${\mathbf{x}}_{t}$ chosen by ${\mathcal{A}}^{\text{action}}$
       Let $f_{t}$ be the function chosen by the adversary
       Query the oracle at the point ${\mathbf{x}}_{t}$ to get ${\mathbf{o}}_{t}$
       Let $q_{t}:={\mathbf{y}}\mapsto\langle{\mathbf{o}}_{t},{\mathbf{y}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}_{t}\|^{2}$
       for $i$ starting from 1, while ${\mathcal{A}}^{\text{query}}$ is not terminated for this time-step do
             Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
             Return $\nabla q_{t}({\mathbf{y}}_{t,i})={\mathbf{o}}_{t}+\mu\langle{\mathbf{y}}_{t,i},{\mathbf{y}}_{t,i}-{\mathbf{x}}_{t}\rangle$ as the output of the query oracle to ${\mathcal{A}}$
       end for
end for
Algorithm 1 Full information to semi-bandit - $\operatorname{FTS}({\mathcal{A}})$
Theorem 3.
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ be a convex set, let $\mu\geq 0$ and let ${\mathcal{A}}$ be algorithm for online optimization with full information first order feedback. If ${\mathbf{F}}$ is a $\mu$-strongly convex function class over ${\mathcal{K}}$ , then we have
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
where ${\mathcal{A}}^{\prime}=\operatorname{FTS}({\mathcal{A}})$ is Algorithm 1. In particular, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
See Appendix C for proof. Note that when the base algorithm ${\mathcal{A}}$ is semi-bandit, we have $\operatorname{FTS}({\mathcal{A}})={\mathcal{A}}$ and the above theorem becomes trivial.
Remark 2.
In Zhang et al., 2018a , Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback. Algorithm 3, i.e., “Improved Ader”, is the result of applying the $\operatorname{FTS}$ meta-algorithm to to Ader.
7 Lipschitz to non-Lipschitz
It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in Flaxman et al., (2005) and Lemma 1). Here we use a shrinking method described in Pedramfar et al., (2023). We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}^{\operatorname{aff}({\mathcal{K}})}({\mathbf{c}})\subseteq{\mathcal{K}}$. Then, for any shrinking parameter $0\leq\alpha<r$, we define $\hat{{\mathcal{K}}}_{\alpha}:=(1-\frac{\alpha}{r}){\mathcal{K}}+\frac{\alpha}{r}{\mathbf{c}}$. We define ${\mathbf{F}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to function class $\{f|_{\hat{{\mathcal{K}}}_{\alpha}}\mid f\in{\mathbf{F}}\}$. Given an adversary $\operatorname{Adv}$, we define $\operatorname{Adv}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the adversary constructed by restricting the output of $\operatorname{Adv}$ to the set $\hat{{\mathcal{K}}}_{\alpha}$. Recall that the the domain ${\mathcal{K}}$ of the adversary is known to the agent ${\mathcal{A}}$. For an online algorithm ${\mathcal{A}}$, we define ${\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the online algorithm resulting from restricting the domain of ${\mathcal{A}}$ to $\hat{{\mathcal{K}}}_{\alpha}$.
Theorem 4.
Let ${\mathcal{A}}$ be an online algorithm and ${\mathbf{F}}$ be a convex function class bounded by $M_{0}$. Also let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and $\hat{{\mathcal{U}}}:=(1-\frac{\alpha}{r}){\mathcal{U}}+\frac{\alpha}{r}{\mathbf{c}}$. Then, for any adversary $\operatorname{Adv}$, we have
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}.$
See Appendix D for proof.
8 Deterministic feedback to stochastic feedback
In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. In particular, there is no need for any variance reduction method, such as momentum Mokhtari et al., (2020); Chen et al., (2018); Xie et al., (2020), as long as we know that the algorithm is designed for a fully adaptive adversary.
Theorem 5.
Let ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ be a convex set, let $\mu\geq 0$ and let ${\mathcal{A}}$ be an algorithm for online optimization with semi-bandit feedback. If ${\mathbf{F}}$ is an $M_{1}$-Lipschitz $\mu$-strongly convex function class over ${\mathcal{K}}$ and $B_{1}\geq M_{1}$, then we have
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
See Appendix E for proof. Theorem 6.5 in Hazan et al., (2016) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
9 First order feedback to zeroth order feedback
When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient. The “smoothing trick” is a classical result by Flaxman et al., (2005) which involves averaging through spherical sampling around a given point. Here we use a variant that was introduced in Pedramfar et al., (2023) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm ${\mathcal{A}}$
Pass $\hat{{\mathcal{K}}}_{\alpha}$ as the domain to ${\mathcal{A}}$
$k\leftarrow\operatorname{dim}({\mathcal{L}}_{0})$
for $t=1,2,\dots,T$ do
       ${\mathbf{x}}_{t}\leftarrow$ the action chosen by ${\mathcal{A}}$
       Play ${\mathbf{x}}_{t}$
       Let $f_{t}$ be the function chosen by the adversary
       for $i$ starting from 1, while ${\mathcal{A}}^{\text{query}}$ is not terminated for this time-step do
             Sample ${\mathbf{v}}_{t,i}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
             Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
             Query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$ to get $o_{t,i}$
             Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
       end for
end for
Algorithm 2 First order to zeroth order - $\operatorname{FOTZO}({\mathcal{A}})$
For a function $f:{\mathcal{K}}\to{\mathbb{R}}$ defined on a convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ and a smoothing parameter $0<\delta\leq\alpha$, its $\delta$-smoothed version $\hat{f}_{\delta}:\hat{{\mathcal{K}}}_{\alpha}\to{\mathbb{R}}$ is given as
$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}^{\operatorname{aff}({\mathcal{K}})}({\mathbf{x}})}[f({\mathbf{z}})]={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}^{{\mathcal{L}}_{0}}({\mathbf{0}})}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
where ${\mathcal{L}}_{0}=\operatorname{aff}({\mathcal{K}})-{\mathbf{x}}$, for any ${\mathbf{x}}\in{\mathcal{K}}$ is the linear space that is a translation of the affine hull of ${\mathcal{K}}$ and ${\mathbf{v}}$ is sampled uniformly at random from the $k=\operatorname{dim}({\mathcal{L}}_{0})$-dimensional ball ${\mathbb{B}}_{1}^{{\mathcal{L}}_{0}}({\mathbf{0}})$. Thus, the function value $\hat{f}_{\delta}({\mathbf{x}})$ is obtained by “averaging” $f$ over a sliced ball of radius $\delta$ around ${\mathbf{x}}$. For a function class ${\mathbf{F}}$ over ${\mathcal{K}}$, we use $\hat{{\mathbf{F}}}_{\delta}$ to denote $\{\hat{f}_{\delta}\mid f\in{\mathbf{F}}\}$. We will drop the subscripts $\alpha$ and $\delta$ when there is no ambiguity.
Theorem 6.
Let ${\mathbf{F}}$ be a function class over a convex set ${\mathcal{K}}$ and choose ${\mathbf{c}}$ and $r$ as described above and let $\delta\leq\alpha<r$. Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\alpha}{r}){\mathcal{U}}+\frac{\alpha}{r}{\mathbf{c}}$. Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback. Then, if ${\mathcal{A}}^{\prime}=\operatorname{FOTZO}({\mathcal{A}})$ where $\operatorname{FOTZO}$ is described by Algorithm 2, we have:
  • if ${\mathbf{F}}$ is convex and bounded by $M_{0}$, we have
    ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T.$
  • if ${\mathbf{F}}$ is $M_{1}$-Lipschitz (but not necessarily convex), then we have
    ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T.$
Theorem 7.
Under the assumptions of Theorem 6, if we further assume that ${\mathcal{A}}$ is semi-bandit, then the same regret bounds hold with ${\mathcal{A}}^{\prime}=\operatorname{STB}({\mathcal{A}})$, where $\operatorname{STB}$ is described by Algorithm 3.
See Appendix F for proof.
Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm ${\mathcal{A}}$
Pass $\hat{{\mathcal{K}}}_{\alpha}$ as the domain to ${\mathcal{A}}$
$k\leftarrow\operatorname{dim}({\mathcal{L}}_{0})$
for $t=1,2,\dots,T$ do
       Sample ${\mathbf{v}}_{t}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
       ${\mathbf{x}}_{t}\leftarrow$ the action chosen by ${\mathcal{A}}$
       Play ${\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t}$
       Let $f_{t}$ be the function chosen by the adversary
       Let $o_{t}$ be the output of the value oracle
       Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
end for
Algorithm 3 Semi-bandit to bandit - $\operatorname{STB}({\mathcal{A}})$
Remark 3.
While $\operatorname{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\operatorname{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm. However, we note that even when ${\mathcal{A}}$ is semi-bandit, $\operatorname{FOTZO}({\mathcal{A}})$ is not a bandit algorithm. Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action. This is because $\operatorname{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm ${\mathcal{A}}$, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
Remark 4.
We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough. However, we have included the $\operatorname{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
Recall that OGD algorithm Zinkevich, (2003) is a deterministic algorithm designed for deterministic semi-bandit feedback. Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result. This generalizes the result of Flaxman et al., (2005) to allow the feedback of the value oracle to be stochastic.
Corollary 3.
We have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(DB_{0}T^{3/4}\right).$
As mentioned in Remark 2, the Improved Ader algorithm of Zhang et al., 2018a is a deterministic algorithm designed for deterministic semi-bandit feedback. Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result. This matches the SOTA result of Zhao et al., (2021) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback.
Corollary 4.
Let IA denote the algorithm mentioned above. Also let ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$. Then we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{IA})}({\mathbf{u}})[1,T]$ $=O\left(DB_{0}\left(1+\frac{P_{T}({\mathbf{u}})}{D}\right)^{1/2}T^{3/4}\right).$
Similarly, the SO-OGD algorithm Garber and Kretzu, (2022) is a deterministic algorithm designed for deterministic semi-bandit feedback with obtains a weakly adaptive regret bound of $O(DM_{1}T^{1/2})$ (Theorem 14 in Garber and Kretzu, (2022)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result. This matches the result of Theorem 15 in Garber and Kretzu, (2022) and generalizes it to allow for stochastic feedback.
Corollary 5.
Let SO-OGD be the algorithm mentioned above. Then we have
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{SO-OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(DB_{0}T^{3/4}\right).$
10 Acknowledgements
The authors want to acknowledge the useful feedback on the paper and its ideas provided by Yuanyu Wan.
Appendix A Proof of Theorem 1
Proof.
Note that $\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}})\subseteq\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})$ which directly implies that
${\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}\geq{\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}^{{\mathcal{A}}}.$
To prove the inequality in the other direction, let ${\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})$. The game between $({\mathcal{A}},{\mathcal{B}})$ is a deterministic game and only a single sequence of functions $(f_{1},\cdots,f_{T})$ could be generated by ${\mathcal{B}}$, i.e., there is a unique ${\mathcal{B}}^{\prime}:=(f_{1},\cdots,f_{T})\in\operatorname{Adv}^{\text{o}}_{i}({\mathbf{F}})$ such that the game played by $({\mathcal{A}},{\mathcal{B}})$ is identical to the game played by $({\mathcal{A}},{\mathcal{B}}^{\prime})$. Therefore, we have
${\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}^{\prime}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}^{\prime}}={\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}^{{\mathcal{A}}}.\qed$
Appendix B Proof of Theorem 2
Proof.
Regret:
Let ${\mathbf{o}}_{t}$ denote the output of the subgradient query oracle at time-step $t$. For any realization ${\mathcal{B}}=({\mathcal{B}}_{1},\cdots,{\mathcal{B}}_{T})\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})$, we define ${\mathcal{B}}^{\prime}_{t}({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})$ to be the tuple $(q_{t},\nabla)$ where
$q_{t}:={\mathbf{y}}\mapsto\langle{\mathbf{o}}_{t},{\mathbf{y}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}_{t}\|^{2}\in{\mathbf{Q}}_{\mu}.$
We also define ${\mathcal{B}}^{\prime}:=({\mathcal{B}}^{\prime}_{1},\cdots,{\mathcal{B}}^{\prime}_{T})$. Note that each ${\mathcal{B}}^{\prime}_{t}$ is a function of ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$ and therefore it belongs to $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})$. Since the algorithm uses semi-bandit feedback, the sequence of random vectors $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ chosen by ${\mathcal{A}}$ is identical between the game with ${\mathcal{B}}$ and ${\mathcal{B}}^{\prime}$. Therefore, for any ${\mathbf{y}}\in{\mathcal{K}}$, we have
$q_{t}({\mathbf{x}}_{t})-q_{t}({\mathbf{y}})=\langle{\mathbf{o}}_{t},{\mathbf{x}}_{t}-{\mathbf{y}}\rangle-\frac{\mu}{2}\|{\mathbf{x}}_{t}-{\mathbf{y}}\|^{2}\geq f_{t}({\mathbf{x}}_{t})-f_{t}({\mathbf{y}}),$
where the last inequality follows from fact that ${\mathbf{o}}_{t}$ is a subgradient of $f_{t}$ at ${\mathbf{x}}_{t}$. Therefore, we have
$\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\leq\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right),$ (1)
which implies that
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}{\mathbb{E}}\left[\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right]$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}{\mathbb{E}}\left[\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right)\right]\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
To prove the second claim, we note that ${\mathbf{F}}\supseteq{\mathbf{F}}_{\mu}$ implies that $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})\supseteq\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})$. Therefore
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}_{{\mathcal{B}},\nabla^{*}}^{{\mathcal{A}}}\geq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}{\mathcal{R}}_{{\mathcal{B}},\nabla^{*}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
High probability regret bound:
Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 1, we see that
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}=\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\leq\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right)={\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}},$
which immediately implies that
(${\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
Hence $h$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}))$ as well.
Moreover, if ${\mathbf{F}}\supseteq{\mathbf{F}}_{\mu}$, then $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})\supseteq\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})$ and therefore, for any function $h:[0,1]\to[0,\infty)$, we have
$\inf_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)$ $\leq\inf_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right).$
Hence any high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}))$ is one for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$ as well. ∎
Appendix C Proof of Theorem 3
Proof.
First we note that for any adversary $\operatorname{Adv}$ over ${\mathbf{Q}}_{\mu}$ with deterministic first order feedback, under the algorithms ${\mathcal{A}}^{\prime}$, we have $q_{t}=f_{t}$. Hence ${\mathcal{A}}^{\prime}$ and ${\mathcal{A}}$ return the same sequence of points and we have
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}^{\prime}}$ $={\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}\qquad\text{and}\qquad\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}^{\prime}}=\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}}$
Therefore, using Theorem 2 and the fact that ${\mathbf{F}}_{\mu}\subseteq{\mathbf{Q}}_{\mu}$, we see that
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}$ $\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}^{\prime}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}},$
and similarly
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}.$
As a special case, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then we may use the same argument in the proof of Theorem 2 to see that
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
and
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
which completes the proof. ∎
Appendix D Proof of Theorem 4
Proof.
First note that for any ${\mathcal{B}}\in\operatorname{Adv}$, the game played between $({\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}},{\mathcal{B}})$ is identical to the game played between $({\mathcal{A}},{\mathcal{B}}|_{\hat{{\mathcal{K}}}_{\alpha}})$. Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$. We have
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}-{\mathcal{R}}_{{\mathcal{B}}|_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}$ $={\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]-{\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})\right]$
$={\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right].$
We have
$\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$
$=\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$\leq\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}f_{t}({\mathbf{c}})\right)$ (Convexity)
$\leq\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}M_{0}\right)$
$=\frac{\alpha M_{0}T}{r}+\left(1-\frac{\alpha}{r}\right)\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
$\leq\frac{\alpha M_{0}T}{r}+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t}).$
Putting these together, we see that
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})-{\mathcal{R}}_{{\mathcal{B}}|_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})$ $\leq\frac{\alpha M_{0}T}{r}.$ (2)
Therefore, we have
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathcal{R}}_{{\mathcal{B}}|_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}={\mathcal{R}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}.$
The proof of the claim for high probability regret bounds follows from Equation 2 and the same argument used in the last part of the proof of Theorem 2. ∎
Appendix E Proof of Theorem 5
Proof.
Let $\Omega^{\mathcal{Q}}=\Omega^{\mathcal{Q}}_{1}\times\cdots\times\Omega^{\mathcal{Q}}_{T}$ capture all sources of randomness in the query oracles of $\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})$, i.e., for any choice of $\theta\in\Omega^{\mathcal{Q}}$, the query oracle is deterministic. Hence for any $\theta\in\Omega^{\mathcal{Q}}$ and realized adversary ${\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})$, we may consider ${\mathcal{B}}_{\theta}$ as an object similar to an adversary with a deterministic oracle. However, note that ${\mathcal{B}}_{\theta}$ does not satisfy the unbiasedness condition of the oracle, i.e., the returned value of the oracle is not necessarily the gradient of the function at that point. Recall that $h_{t}$ denotes the history and ${\mathcal{B}}_{t}$ maps a tuple $(h_{t},{\mathbf{x}}_{t})$ to a tuple of $f_{t}$ and a stochastic query oracle for $f_{t}$. We will use ${\mathbb{E}}_{\Omega^{\mathcal{Q}}}$ to denote the expectation with respect to the randomness of query oracle and ${\mathbb{E}}_{\Omega^{\mathcal{Q}}_{t}}[\cdot]:={\mathbb{E}}_{\Omega^{\mathcal{Q}}}[\cdot|f_{t},{\mathbf{x}}_{t}]$ to denote the expectation conditioned on the function and the action of the agent. Similarly, let ${\mathbb{E}}_{\Omega^{\mathcal{A}}}$ denote the expectation with respect to the randomness of the agent. Let ${\mathbf{o}}_{t}$ be the random variable denoting the output of the subgradient query oracle at time-step $t$ and let
$\bar{{\mathbf{o}}}_{t}:={\mathbb{E}}[{\mathbf{o}}_{t}\mid f_{t},{\mathbf{x}}_{t}]={\mathbb{E}}_{\Omega^{\mathcal{Q}}_{t}}[{\mathbf{o}}_{t}]$
be a subgradient of $f_{t}$ at ${\mathbf{x}}_{t}$.
For any realization ${\mathcal{B}}=(f_{1},\cdots,f_{T})\in\operatorname{Adv}^{\text{o}}({\mathbf{F}})$ and any $\theta\in\Omega^{\mathcal{Q}}$, we define ${\mathcal{B}}^{\prime}_{\theta,t}({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t})$ to be the pair $(q_{t},\nabla)$ where
$q_{t}:={\mathbf{y}}\mapsto\langle{\mathbf{o}}_{t},{\mathbf{y}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}_{t}\|^{2}\in{\mathbf{Q}}_{\mu}.$
We also define ${\mathcal{B}}^{\prime}_{\theta}:=({\mathcal{B}}^{\prime}_{\theta,1},\cdots,{\mathcal{B}}^{\prime}_{\theta,T})$. Note that a specific choice of $\theta$ is necessary to make sure that the function returned by ${\mathcal{B}}^{\prime}_{\theta,t}$ is a deterministic function of ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$ and not a random variable and therefore ${\mathcal{B}}^{\prime}_{\theta}$ belongs to $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])$.
Since the algorithm uses (semi-)bandit feedback, given a specific value of $\theta$, the sequence of random vectors $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ chosen by ${\mathcal{A}}$ is identical between the game with ${\mathcal{B}}_{\theta}$ and ${\mathcal{B}}^{\prime}_{\theta}$. Therefore, for any ${\mathbf{u}}_{t}\in{\mathcal{K}}$, we have
$f_{t}({\mathbf{x}}_{t})-f_{t}({\mathbf{u}}_{t})$ $\leq\langle\bar{{\mathbf{o}}}_{t},{\mathbf{x}}_{t}-{\mathbf{u}}_{t}\rangle-\frac{\mu}{2}\|{\mathbf{x}}_{t}-{\mathbf{u}}_{t}\|^{2}$
$=\langle{\mathbb{E}}\left[{\mathbf{o}}_{t}\mid h_{t},{\mathbf{x}}_{t}\right],{\mathbf{x}}_{t}-{\mathbf{u}}_{t}\rangle-\frac{\mu}{2}\|{\mathbf{x}}_{t}-{\mathbf{u}}_{t}\|^{2}$
$={\mathbb{E}}\left[\langle{\mathbf{o}}_{t},{\mathbf{x}}_{t}-{\mathbf{u}}_{t}\rangle-\frac{\mu}{2}\|{\mathbf{x}}_{t}-{\mathbf{u}}_{t}\|^{2}\mid f_{t},{\mathbf{x}}_{t}\right]$
$={\mathbb{E}}\left[q_{t}({\mathbf{x}}_{t})-q_{t}({\mathbf{u}}_{t})\mid h_{t},{\mathbf{x}}_{t}\right]$
$={\mathbb{E}}_{\Omega^{\mathcal{Q}}_{t}}\left[q_{t}({\mathbf{x}}_{t})-q_{t}({\mathbf{u}}_{t})\right]$
where the first inequality follows from the fact that $f_{t}$ is $\mu$-strongly convex and $\bar{{\mathbf{o}}}_{t}$ is a subgradient of $f_{t}$ at ${\mathbf{x}}_{t}$. Therefore we have
${\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]\leq{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\sum_{t=a}^{b}{\mathbb{E}}_{\Omega^{\mathcal{Q}}_{t}}\left[q_{t}({\mathbf{x}}_{t})-q_{t}({\mathbf{u}}_{t})\right]\right]={\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-q_{t}({\mathbf{u}}_{t})\right].$
Since ${\mathcal{B}}$ is oblivious, the sequence $(f_{1},\cdots,f_{T})$ is not affected by the randomness of query oracles. Therefore we have
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}$ $={\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]$
$=\max_{{\mathbf{u}}\in{\mathcal{U}}}{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]$
$\leq\max_{{\mathbf{u}}\in{\mathcal{U}}}{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right]$
$\leq{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right)\right]={\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\right],$
where the second inequality follows from Jensen’s inequality.
Regret:
By taking the expectation with respect to $\omega\in\Omega^{\mathcal{A}}$, we see that
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}={\mathbb{E}}_{\Omega^{\mathcal{A}}}\left[{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\right]\leq{\mathbb{E}}_{\Omega^{\mathcal{A}}}\left[{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\right]\right]={\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}\right]\leq\sup_{\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}$
Hence we have
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
High probability regret bound:
By taking the supremum over ${\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})$, we see that
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}_{\omega}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\ \right]$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}_{\omega}}$
Therefore we may use the same argument in the proof of Theorem 2 to see that
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
Appendix F Proof of Theorems 6 and 7
First we prove the following lemma which is a simple generalization of Observation 3 in Flaxman et al., (2005) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
Lemma 1.
For any ${\mathbf{x}}\in\hat{{\mathcal{K}}}_{\alpha}$ and ${\mathbf{y}}\in{\mathcal{K}}$ and any convex function $f:{\mathcal{K}}\to[-M_{0},M_{0}]$, we have
$|f({\mathbf{x}})-f({\mathbf{y}})|\leq\frac{2M_{0}}{\alpha}\|{\mathbf{x}}-{\mathbf{y}}\|$
Proof.
Let ${\mathbf{y}}={\mathbf{x}}+\Delta$. If $\|\Delta\|\geq\alpha$, then we have
$|f({\mathbf{x}})-f({\mathbf{y}})|\leq 2M_{0}\leq\frac{2M_{0}}{\alpha}\|{\mathbf{x}}-{\mathbf{y}}\|.$
Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$. According to Lemma 7 in Pedramfar et al., (2023), we have ${\mathbf{z}}\in{\mathcal{K}}$. We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$. Therefore, since $f$ is convex and bounded by $M_{0}$, we have
$f({\mathbf{y}})$ $\leq\frac{\|\Delta\|}{\alpha}f({\mathbf{z}})+\left(1-\frac{\|\Delta\|}{\alpha}\right)f({\mathbf{x}})=f({\mathbf{x}})+\frac{\|\Delta\|}{\alpha}(f({\mathbf{z}})-f({\mathbf{x}}))\leq f({\mathbf{x}})+\frac{2M_{0}}{\alpha}\|{\mathbf{x}}-{\mathbf{y}}\|.\qed$
We start with a proof of Theorem 7 and then adapt the proof to work for Theorem 6.
Proof of Theorem 7.
Note that any realized adversary ${\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})$ may be represented as a sequence of functions $(f_{1},\cdots,f_{T})$ and a corresponding sequence of query oracles $({\mathcal{Q}}_{1},\cdots,{\mathcal{Q}}_{T})$. For such realized adversary ${\mathcal{B}}$, we define $\hat{B}$ to be the realized adversary corresponding to $(\hat{f}_{1},\cdots,\hat{f}_{T})$ with the stochastic gradient oracles
$\hat{{\mathcal{Q}}}_{t}({\mathbf{x}}):=\frac{k}{\delta}{\mathcal{Q}}_{t}({\mathbf{x}}+\delta{\mathbf{v}}){\mathbf{v}},$ (3)
where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in Pedramfar et al., (2023), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.  [^3] When using a spherical estimator, it was shown in Flaxman et al., (2005) that $\hat{f}$ is differentiable even when $f$ is not. When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in Pedramfar et al., (2023). However, their proof is based on the proof for the spherical case and therefore the results carry forward to show that $\hat{f}$ is differentiable. Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})$. Using Equation 3 and the definition of the Algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the sequence of actions $({\mathbf{x}}_{1}+\delta{\mathbf{v}}_{1},\cdots,{\mathbf{x}}_{T}+\delta{\mathbf{v}}_{T})$ in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$. We have
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $={\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]-{\mathbb{E}}\left[\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})\right]$
$={\mathbb{E}}\left[\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})\right)+\left(\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right].$ (4)
First we consider the convex case. By definition, $\hat{f_{t}}({\mathbf{x}}_{t})$ is an average of $f_{t}$ at points that are within $\delta$ distance of ${\mathbf{x}}$. Therefore, according to Lemma 1 we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\frac{2M_{0}\delta}{\alpha}$. By using Lemma 1 again for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
$|f_{t}({\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq|f_{t}({\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})-f_{t}({\mathbf{x}}_{t})|+|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\frac{4M_{0}\delta}{\alpha}.$ (5)
On the other hand, we have
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 1)
$=\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}f_{t}({\mathbf{c}})\right)$ (Convexity)
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}M_{0}\right)$
$=\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\left(1-\frac{\alpha}{r}\right)\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
$\leq\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t}).$
Putting these together, we see that
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $\leq\frac{4M_{0}\delta}{\alpha}T+\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T=\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T.$
Therefore, we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}{\mathcal{R}}_{{\mathcal{B}},{\mathcal{Q}}}^{{\mathcal{A}}^{\prime}}$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T$
$\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}+\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T.$
To prove the result in the Lipschitz case, we note that, according to Lemma 3 in Pedramfar et al., (2023), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
$|f_{t}({\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq|f_{t}({\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})-f_{t}({\mathbf{x}}_{t})|+|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq 2\delta M_{1}.$ (6)
On the other hand, we have
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in Pedramfar et al., (2023))
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\alpha}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\alpha M_{1}D}{r}\right)$ (Lipschitz)
$=\left(1+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
Therefore, using Equation 4, we see that
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $\leq 2\delta M_{1}T+\left(1+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T=\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T.$
Therefore, we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T$
$\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T.$
Proof of Theorem 6.
The proof of the bounds for this case is similar to the previous case. As before, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows from the description of Algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$. We have
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $={\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]-{\mathbb{E}}\left[\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})\right]$
$={\mathbb{E}}\left[\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})\right)+\left(\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right].$ (7)
To obtain the same bound as before, we note that in the convex case, instead of Inequality 5, we have
$|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\frac{2M_{0}\delta}{\alpha}<\frac{4M_{0}\delta}{\alpha},$ (8)
and, in the Lipschitz case, instead of Inequality F, we have
$|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\delta M_{1}<2\delta M_{1}.$
The rest of the proof follows verbatim. ∎

### v3

@@
-A Generalized Approach to Online Convex Optimization
+A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization
 Abstract
-In this paper, we analyze the problem of online convex optimization in different settings.
+In this paper, we analyze the problem of online convex optimization in different settings, including different feedback types (full-information/semi-bandit/bandit/etc) in either stochastic or non-stochastic setting and different notions of regret (static adversarial regret/dynamic regret/adaptive regret).
+This is done through a framework which allows us to systematically propose and analyze meta-algorithms for the various settings described above.
 We show that any algorithm for online linear optimization with fully adaptive adversaries is an algorithm for online convex optimization.
@@
 We use this to describe general meta-algorithms to convert first order algorithms to zeroth order algorithms with comparable regret bounds.
-Our framework allows us to analyze online optimization in various settings, such full-information feedback, bandit feedback, stochastic regret, adversarial regret and various forms of non-stationary regret, and recovers several results in the literature with a simplified proof technique.
+Our framework allows us to analyze online optimization in various settings, recovers several results in the literature with a simplified proof technique, and provides new results.
 1 Introduction
 1.1 Overview
-Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information Shalev-Shwartz, (2012); Hazan et al., (2016).
+Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information (?, ?).
 This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$.
-The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, the information exchange through a query oracle.
-In this paper, we present a comprehensive approach to solving online convex optimization problems, addressing scenarios where the adversary can choose a function from the class of $\mu$–strongly convex functions (or convex functions for $\mu=0$) at each time step.
+The dynamics of this game involve the agent’s decision-making process, the adversary’s strategy, and the information exchange through a query oracle.
+In this paper, we present a comprehensive approach to solve online convex optimization problems, addressing scenarios where the adversary can choose a function from the class of $\mu$–strongly convex functions (or convex functions for $\mu=0$) at each time step.
 Our approach encompasses different feedback types, including bandit, semi-bandit and full-information feedback in each iteration.
@@
 Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries.
-Finally, we introduce a meta algorithm, based on the variant of classical result of Flaxman et al., (2005), that converts algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit feedback.
+Finally, we introduce meta algorithms, based on the variant of classical result of (?, ?), that convert algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit or zeroth-order feedback.
+In addition to recovering many results in the literature with simplified proofs, we give new static regret guarantees for online strongly convex optimization and adaptive regret guarantees for convex optimization with deterministic zeroth order feedback.
 Figure 1: Summary of the main results
@@
 After that, SB is used to denotes semi-bandit, B to denote bandit and FI to denote full information feedback.
-Finally, the last D or S is used to denotes deterministic or stochastic algorithm.
-1.2 Related works
-The term “online convex optimization” was first defined in Zinkevich, (2003), where Online Gradient Descent (OGD) was proposed, while this setting was first introduced in Gordon, (1999).
-The Follow-The-Perturbed-Leader (FTPL) for online linear optimization was introduced by Kalai and Vempala, (2005), inspired by Follow-The-Leader algorithm originally suggested by Hannan, (1957).
-Follow-The-Regularized-Leader (FTRL) in the context of online convex optimization was coined in Shalev-Shwartz and Singer, (2007); Shalev-Shwartz, (2007).
-The equivalence of FTRL and Online Mirror Descent (OMD) was observed by Hazan and Kale, (2010).
-Non-stationary measures of regret have been studied in  Zinkevich, (2003); Hazan and Seshadhri, (2009); Besbes et al., (2015); Daniely et al., (2015); Zhang et al., 2018b ; Zhang et al., 2018a ; Zhao et al., (2020, 2021); Lu et al., (2023); Wang et al., (2024); Garber and Kretzu, (2022) and many other papers.
-We refer the reader to surveys Shalev-Shwartz, (2012) and Hazan et al., (2016) for more details.
+Finally, the last D or S, if exists, is used to denotes deterministic or stochastic algorithm.
 1.2 Contributions
 The key results of this paper are encapsulated in Figure 1, where each arrow represents a procedure or meta-algorithm that transforms an algorithm from one setting to another.
-The primary contribution is two-fold:
-  • we provide a framework that simplifies analysis of meta-algorithms in online learning; and
+The primary contribution is three-fold:
+  • (I)
+    We provide a framework that simplifies analysis of meta-algorithms in online learning.
+  • (II)
     Using this framework, we formulate several meta-algorithms that facilitate the conversion of algorithms across various settings.
-This framework not only simplifies proof analysis for numerous existing problems in the field (see Remark 2 and Corollaries 34 and 5) but also gives a tool for easier extension of multiple results in the literature.
+  • (III)
+    In particular, we convert several insights in the literature into precise theorems, for example the idea of using 1-point gradient estimator in online optimization is formalized in Meta-algorithms 2 and 3 and Theorems 6 and 7.
+Similarly, the insight that “knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere” (See (?)) is formalized in Meta-algorithm 4 and Theorem 8.
+This framework not only simplifies proof analysis for numerous existing problems in the field (see Section 11) but also provides a tool for easier extension of multiple results in the literature.
 In the following, we go over some of the main meta-algorithms that we discuss in this paper.
-The first key result in this paper is Theorem 2.
+(i) The first key result in this paper is Theorem 2.
 This result shows that any algorithm for online linear optimization with stochastic subgradient feedback results in a comparable regret with if the function class is replaced by a class of convex functions.
 This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback.
-We note that this result recovers the result for online convex game in Abernethy et al., (2008), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
+We note that this result recovers the result for online convex game in (?), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
 While the above result is the key for going from online linear to online convex optimization, the main issue for its applicability is that it requires semi-bandit feedback.
 In the remaining results, we provide meta-algorithms to increase the applicability of this result.
-The second key result, given in Theorem 3, is that for any full-information algorithm for online convex optimization with fully adaptive adversary, there is a corresponding algorithm with regret bound of the same order which only requires semi-bandit feedback.
+(ii) The second key result, given in Theorem 3, is that for any full-information algorithm for online convex optimization with fully adaptive adversary, there is a corresponding algorithm with regret bound of the same order which only requires semi-bandit feedback.
 This allows us to get results for semi-bandit feedback from that of full information.
@@
 This combination has been marked as Theorem 3 in Figure 1.
-We note that the above results explained in Theorems 14 assume that the feedback oracle is deterministic.
+(iii) We note that the above results assume that the feedback oracle is deterministic.
 In the next key result, given in Theorem 5, we relax this assumption.
@@
 1.
-This result gives Theorem 6.5 in Hazan et al., (2016) as a special case.
+This result gives Theorem 6.5 in (?) as a special case.
 Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
-Finally, we note that the above discussion was mostly focused on full information and semi-bandit feedback settings.
-However, in many practical setups, zeroth order or bandit feedback is important.
-In the final key result, given in Theorems 6-7, we use the smoothing trick Flaxman et al., (2005); Pedramfar et al., (2023), in order to estimate the gradient from the value oracle.
+(iv) The above discussion was mostly focused on full information and semi-bandit feedback settings.
+However, in many practical setups, we only have access to zeroth order or bandit feedback.
+In the next key result, given in Theorems 6-7, we use the smoothing trick (?, ?, ?), in order to estimate the gradient from the value oracle.
 Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively.
 These results apply to any optimization problem, not necessarily convex.
-This approach recovers the results in Flaxman et al., (2005) with the base algorithm of OGD, while extends to say that the approach can be applied directly on any base algorithm.
-In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback Zhao et al., (2021); Garber and Kretzu, (2022) with a simplified proof.
+This approach recovers the results in (?) with the base algorithm of Online Gradient Descent ($\mathtt{OGD}$).
+In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback (?, ?, ?) with a simplified proof.
+(v) In the above results, we assumed that we only have access to an unbiased estimate of the zeroth order (or bandit) feedback.
+In (?), the setting of multi-point bandit feedback is introduced.
+Even though their setting and notion of regret is different from our work, one insight is relevant.
+Their analysis suggests that, in some cases, if we have access to exact values of the functions, knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere.
+We formalize this insight by designing a meta-algorithm based on the two-point gradient estimator (?, ?) that allows us to convert algorithms for stochastic first order feedback into algorithms for deterministic zeroth order feedback with the same order of regret bound (Theorem 8).
+Note that this result holds for convex and non-convex optimization as it only relies on the function class being Lipschitz.
+This result is shown to recover the result for dynamic regret for Improved Ader (?), while achieving two new results for adaptive regret for convex function of $O(\sqrt{T})$ and static regret for strongly convex function of $O(\log T)$.
 Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings.
@@
 This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results.
-For instance, results in Abernethy et al., (2008); Hazan et al., (2016); Flaxman et al., (2005); Zhao et al., (2021) as mentioned above can be demonstrated as corollaries of these more generalized results.
+For instance, results in (?, ?, ?, ?, ?, ?) as mentioned above can be demonstrated as corollaries of these more generalized results.
+Further, we provide new results, which shows that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$.
+1.3 Related works
+The term “online convex optimization” was first defined in (?), where Online Gradient Descent ($\mathtt{OGD}$) was proposed, while this setting was first introduced in (?).
+The Follow-The-Perturbed-Leader ($\mathtt{FTPL}$) for online linear optimization was introduced by (?), inspired by Follow-The-Leader algorithm originally suggested by (?).
+Follow-The-Regularized-Leader ($\mathtt{FTRL}$) in the context of online convex optimization was coined in (?, ?).
+The equivalence of $\mathtt{FTRL}$ and Online Mirror Descent ($\mathtt{OMD}$) was observed by (?).
+Non-stationary measures of regret have been studied in  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and many other papers.
+We refer the reader to surveys (?) and (?) for more details.
 2 Background and Notations
-For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{K}}$ and $\alpha\in{\mathbb{R}}$.
+For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{D}}$ and $\alpha\in{\mathbb{R}}$.
 The relative interior of ${\mathcal{D}}$ is defined as
 $\operatorname{relint}({\mathcal{D}}):=\{{\mathbf{x}}\in{\mathcal{D}}\mid\exists r>0,{\mathbb{B}}_{r}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{D}})\subseteq{\mathcal{D}}\}.$
+We use ${\mathbb{B}}_{r}({\mathcal{D}})$ to denote the union of balls of radius $r$ centered at ${\mathcal{D}}$.
 A function class is simply a set of real-valued functions.
@@
 Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is called convex if, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $f(\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}})\leq\alpha f({\mathbf{x}})+(1-\alpha)f({\mathbf{y}})$.
-All convex functions are continuous on any point in the relative interior of their domains.
+All convex functions are continuous on any point in the relative interior of their domains, but not necessarily at their relative boundary.
 In this work, we will only focus on continuous functions.
@@
 Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is $\mu$-strongly convex if and only if it has a $\mu$-subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$.
+We use $\nabla^{*}f$ to denote the set of $\mu$-subgradients of $f$.
 3 Problem Setup
@@
 Then, for $1\leq i\leq k_{t}$, the agent chooses a points ${\mathbf{y}}_{t,i}$ and receives the output of the query oracle.
+Here $k_{t}$ denotes the total number of queries made by the agent at time-step $t$, which may or may not be known in advance.
 To be more precise, an agent consists of a tuple $(\Omega^{\mathcal{A}},{\mathcal{A}}^{\text{action}},{\mathcal{A}}^{\text{query}})$, where $\Omega^{\mathcal{A}}$ is a probability space that captures all the randomness of ${\mathcal{A}}$.
@@
 Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$.
-When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret Zinkevich, (2003); Zhang et al., 2018a .
-If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret Besbes et al., (2015).
-Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ Hazan and Seshadhri, (2009).
+When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret (?, ?).
+If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret (?).
+Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ (?).
 We drop $a$, $b$ and ${\mathcal{U}}$ when the statement is independent of their value or their value is clear from the context.
@@
 The following theorem states a result that is well-known in the literature.
-For example, Hazan et al., (2016) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary.
+For example, (?) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary.
 Here we explicitly mention it as a theorem for completion.
@@
 Corollary 1.
-The OGD, FTRL, and OMD algorithms are deterministic and therefore have same regret bound in fully-adaptive setting as the oblivious setting.
+The $\mathtt{OGD}$, $\mathtt{FTRL}$, and $\mathtt{OMD}$ algorithms are deterministic and therefore have same regret bound in fully-adaptive setting as the oblivious setting.
 5 Linear to convex with fully adaptive adversary
@@
 We say ${\mathbf{F}}$ is closed under $\mu$-quadratization if ${\mathbf{F}}\supseteq{\mathbf{F}}_{\mu}$.
-Similarly, for $B>0$, we define ${\mathbf{F}}_{\mu}[B]$ to be the the class of functions $q$ defined above where instead we have ${\mathbf{o}}\in{\mathbb{B}}_{{\mathbf{0}}}(B)$.
+Similarly, for $B>0$, we define ${\mathbf{Q}}_{\mu}[B]$ to be the the class of functions $q$ defined above where instead we have ${\mathbf{o}}\in{\mathbb{B}}_{B}({\mathbf{0}})$.
 Theorem 2.
@@
 See Appendix B for proof.
-Corollary 2 (Theorem 14 in Abernethy et al., (2008)).
+Corollary 2 (Theorem 14 in (?)).
 The min-max regret bound of the problem of online linear optimization with deterministic feedback and online convex optimization with deterministic feedback are equal.
@@
 Specifically, we show that an algorithm that requires full information feedback could be converted to an algorithm that only requires semi-bandit feedback with the same regret bounds.
-The meta-algorithm that does this conversion is described in Algorithm 1.
-Next we show that $\operatorname{FTS}({\mathcal{A}})$ always performs at least as good as ${\mathcal{A}}$ when the oracle is deterministic.
+The meta-algorithm that does this conversion is described in Meta-algorithm 1.
+Next we show that $\mathtt{FTS}({\mathcal{A}})$ always performs at least as good as ${\mathcal{A}}$ when the oracle is deterministic.
 Input :  horizon $T$, algorithm ${\mathcal{A}}=({\mathcal{A}}^{\text{action}},{\mathcal{A}}^{\text{query}})$, strong convexity coefficient $\mu\geq 0$
@@
 end for
-Algorithm 1 Full information to semi-bandit - $\operatorname{FTS}({\mathcal{A}})$
+Meta-algorithm 1 Full information to semi-bandit - $\mathtt{FTS}({\mathcal{A}})$
 Theorem 3.
@@
 ${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
-where ${\mathcal{A}}^{\prime}=\operatorname{FTS}({\mathcal{A}})$ is Algorithm 1. In particular, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then
+where ${\mathcal{A}}^{\prime}=\mathtt{FTS}({\mathcal{A}})$ is Meta-algorithm 1. In particular, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then
 ${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
-See Appendix C for proof. Note that when the base algorithm ${\mathcal{A}}$ is semi-bandit, we have $\operatorname{FTS}({\mathcal{A}})={\mathcal{A}}$ and the above theorem becomes trivial.
-Remark 2.
-In Zhang et al., 2018a , Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback.
-Algorithm 3, i.e., “Improved Ader”, is the result of applying the $\operatorname{FTS}$ meta-algorithm to to Ader.
+See Appendix C for proof. Note that when the base algorithm ${\mathcal{A}}$ is semi-bandit, we have $\mathtt{FTS}({\mathcal{A}})={\mathcal{A}}$ and the above theorem becomes trivial.
 7 Lipschitz to non-Lipschitz
-It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in Flaxman et al., (2005) and Lemma 1).
-Here we use a shrinking method described in Pedramfar et al., (2023).
-We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}^{\operatorname{aff}({\mathcal{K}})}({\mathbf{c}})\subseteq{\mathcal{K}}$.
+It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in (?) and Lemma 1).
+Here we use a shrinking method described in (?).
+We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}({\mathbf{c}})\cap\operatorname{aff}({\mathcal{K}})\subseteq{\mathcal{K}}$.
 Then, for any shrinking parameter $0\leq\alpha<r$, we define $\hat{{\mathcal{K}}}_{\alpha}:=(1-\frac{\alpha}{r}){\mathcal{K}}+\frac{\alpha}{r}{\mathbf{c}}$.
@@
 In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries.
-In particular, there is no need for any variance reduction method, such as momentum Mokhtari et al., (2020); Chen et al., (2018); Xie et al., (2020), as long as we know that the algorithm is designed for a fully adaptive adversary.
+In particular, there is no need for any variance reduction method, such as momentum (?, ?, ?), as long as we know that the algorithm is designed for a fully adaptive adversary.
 Theorem 5.
@@
 If ${\mathbf{F}}$ is an $M_{1}$-Lipschitz $\mu$-strongly convex function class over ${\mathcal{K}}$ and $B_{1}\geq M_{1}$, then we have
-${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
-See Appendix E for proof. Theorem 6.5 in Hazan et al., (2016) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
+${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
+See Appendix E for proof. Theorem 6.5 in (?) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
 9 First order feedback to zeroth order feedback
 When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient.
-The “smoothing trick” is a classical result by Flaxman et al., (2005) which involves averaging through spherical sampling around a given point.
-Here we use a variant that was introduced in Pedramfar et al., (2023) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
+The “smoothing trick” is a classical idea in convex optimization (See (?)) which was first used in online convex optimization in (?).
+This idea involves averaging through spherical sampling around a given point.
+Here we use a variant that was introduced in (?) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
 Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm ${\mathcal{A}}$
@@
 for $t=1,2,\dots,T$ do
-       ${\mathbf{x}}_{t}\leftarrow$ the action chosen by ${\mathcal{A}}$
-       Play ${\mathbf{x}}_{t}$
+       Play ${\mathbf{x}}_{t}$, where ${\mathbf{x}}_{t}$ is the action chosen by ${\mathcal{A}}$
        Let $f_{t}$ be the function chosen by the adversary
@@
              Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
-             Query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$ to get $o_{t,i}$
+             Let $o_{t,i}$ be the output of the query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$
              Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
@@
 end for
-Algorithm 2 First order to zeroth order - $\operatorname{FOTZO}({\mathcal{A}})$
+Meta-algorithm 2 First order to zeroth order - $\mathtt{FOTZO}({\mathcal{A}})$
+Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm ${\mathcal{A}}$
+Pass $\hat{{\mathcal{K}}}_{\alpha}$ as the domain to ${\mathcal{A}}$
+$k\leftarrow\operatorname{dim}({\mathcal{L}}_{0})$
+for $t=1,2,\dots,T$ do
+       Sample ${\mathbf{v}}_{t}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
+       Play ${\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t}$, where ${\mathbf{x}}_{t}$ is the action chosen by ${\mathcal{A}}$
+       Let $f_{t}$ be the function chosen by the adversary
+       Let $o_{t}$ be the output of the value oracle
+       Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
+end for
+Meta-algorithm 3 Semi-bandit to bandit - $\mathtt{STB}({\mathcal{A}})$
 For a function $f:{\mathcal{K}}\to{\mathbb{R}}$ defined on a convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ and a smoothing parameter $0<\delta\leq\alpha$, its $\delta$-smoothed version $\hat{f}_{\delta}:\hat{{\mathcal{K}}}_{\alpha}\to{\mathbb{R}}$ is given as
-$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}^{\operatorname{aff}({\mathcal{K}})}({\mathbf{x}})}[f({\mathbf{z}})]={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}^{{\mathcal{L}}_{0}}({\mathbf{0}})}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
-where ${\mathcal{L}}_{0}=\operatorname{aff}({\mathcal{K}})-{\mathbf{x}}$, for any ${\mathbf{x}}\in{\mathcal{K}}$ is the linear space that is a translation of the affine hull of ${\mathcal{K}}$ and ${\mathbf{v}}$ is sampled uniformly at random from the $k=\operatorname{dim}({\mathcal{L}}_{0})$-dimensional ball ${\mathbb{B}}_{1}^{{\mathcal{L}}_{0}}({\mathbf{0}})$.
+$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{K}})}[f({\mathbf{z}})]$
+$={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
+where ${\mathcal{L}}_{0}=\operatorname{aff}({\mathcal{K}})-{\mathbf{x}}$, for any ${\mathbf{x}}\in{\mathcal{K}}$ is the linear space that is a translation of the affine hull of ${\mathcal{K}}$ and ${\mathbf{v}}$ is sampled uniformly at random from the $k=\operatorname{dim}({\mathcal{L}}_{0})$-dimensional ball ${\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}$.
 Thus, the function value $\hat{f}_{\delta}({\mathbf{x}})$ is obtained by “averaging” $f$ over a sliced ball of radius $\delta$ around ${\mathbf{x}}$.
@@
 Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback.
-Then, if ${\mathcal{A}}^{\prime}=\operatorname{FOTZO}({\mathcal{A}})$ where $\operatorname{FOTZO}$ is described by Algorithm 2, we have:
+Then, if ${\mathcal{A}}^{\prime}=\mathtt{FOTZO}({\mathcal{A}})$ where $\mathtt{FOTZO}$ is described by Meta-algorithm 2, we have:
   • if ${\mathbf{F}}$ is convex and bounded by $M_{0}$, we have
@@
 Theorem 7.
-Under the assumptions of Theorem 6, if we further assume that ${\mathcal{A}}$ is semi-bandit, then the same regret bounds hold with ${\mathcal{A}}^{\prime}=\operatorname{STB}({\mathcal{A}})$, where $\operatorname{STB}$ is described by Algorithm 3.
+Under the assumptions of Theorem 6, if we further assume that ${\mathcal{A}}$ is semi-bandit, then the same regret bounds hold with ${\mathcal{A}}^{\prime}=\mathtt{STB}({\mathcal{A}})$, where $\mathtt{STB}$ is described by Meta-algorithm 3.
 See Appendix F for proof.
-Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm ${\mathcal{A}}$
-Pass $\hat{{\mathcal{K}}}_{\alpha}$ as the domain to ${\mathcal{A}}$
+Remark 2.
+While $\mathtt{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\mathtt{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm.
+However, we note that even when ${\mathcal{A}}$ is semi-bandit, $\mathtt{FOTZO}({\mathcal{A}})$ is not a bandit algorithm.
+Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action.
+This is because $\mathtt{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm ${\mathcal{A}}$, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
+Remark 3.
+We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough.
+However, we have included the $\mathtt{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
+10 First order feedback to deterministic zeroth order feedback
+When we have access to a deterministic zeroth order feedback, we may use the two point gradient estimator (?, ?) to significantly improve the results of the previous section.
+Theorem 8.
+Let ${\mathbf{F}}$ be an $M_{1}$-Lipschitz function class over a convex set ${\mathcal{K}}$ and choose ${\mathbf{c}}$ and $r$ as described above and let $\delta<r$.
+Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\delta}{r}){\mathcal{U}}+\frac{\delta}{r}{\mathbf{c}}$.
+Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback.
+Then, if ${\mathcal{A}}^{\prime}=\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$ where $\mathtt{FOTZO\textrm{-}2P}$ is described by Meta-algorithm 2, we have
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D}{r}\right)\delta M_{1}T.$
+Note that here we may choose $\delta=T^{-1}$ to see that the order of regret of ${\mathcal{A}}^{\prime}$ remains the same as that of ${\mathcal{A}}$. See Appendix G for proof.
+Input :  Shrunk domain $\hat{{\mathcal{K}}}_{\delta}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta<r$, horizon $T$, algorithm ${\mathcal{A}}$
+Pass $\hat{{\mathcal{K}}}_{\delta}$ as the domain to ${\mathcal{A}}$
 $k\leftarrow\operatorname{dim}({\mathcal{L}}_{0})$
 for $t=1,2,\dots,T$ do
-       Sample ${\mathbf{v}}_{t}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
-       ${\mathbf{x}}_{t}\leftarrow$ the action chosen by ${\mathcal{A}}$
-       Play ${\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t}$
+       Play ${\mathbf{x}}_{t}$, where ${\mathbf{x}}_{t}$ is the action chosen by ${\mathcal{A}}$
        Let $f_{t}$ be the function chosen by the adversary
-       Let $o_{t}$ be the output of the value oracle
-       Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
+       for $i$ starting from 1, while ${\mathcal{A}}^{\text{query}}$ is not terminated for this time-step do
+             Sample ${\mathbf{v}}_{t,i}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
+             Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
+             Query the deterministic oracle at the points ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$ and ${\mathbf{y}}_{t,i}-\delta{\mathbf{v}}_{t,i}$
+             Pass $\frac{k}{2\delta}\left(f_{t}({\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i})-f_{t}({\mathbf{y}}_{t,i}-\delta{\mathbf{v}}_{t,i})\right){\mathbf{v}}_{t,i}$ as the oracle output to ${\mathcal{A}}$
        end for
-Algorithm 3 Semi-bandit to bandit - $\operatorname{STB}({\mathcal{A}})$
+end for
+Meta-algorithm 4 First order to zeroth order with Two Point gradient estimator - $\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$
+11 Applications
+In this section, we discuss applications of meta-algorithms discussed above to some specific base algorithms.
+Note that all the results stated here are immediate applications of our framework to existing results and no extra proof is needed.
+Online Gradient Descent: Recall that $\mathtt{OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback that obtains a regret bound of $O\left(M_{1}T^{1/2}\right)$ over $M_{1}$-Lipschitz convex class ${\mathbf{F}}$ and $O\left(M_{1}^{2}\log(T)\right)$ over strongly convex $M_{1}$-Lipschitz convex class ${\mathbf{F}}^{\prime}$ (see (?)).
+Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$ and $\alpha=\delta=(T\log(T))^{-1/2}$, respectively, to see that
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{0}T^{3/4}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]=O\left(B_{0}^{2}T^{1/2}\log(T)^{1/2}\right).$
+The first result generalizes the result of (?) to allow the feedback of the value oracle to be stochastic.
+The second result provides a new optimal algorithm for bandit strongly convex optimization, matching the result of (?).
+However, (?) obtains a regret bound of the same order for the strongly convex case while assuming access to a self-concordant barrier for the domain and also assuming that the functions are smooth.
+We do not require either of these assumptions here.
+On the other hand, using Theorem 8 on $\mathtt{OGD}$, we see that:
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right),$
+where the first result is proven in (?) and the second one is novel.
+Ader and Improved Ader: In (?), Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback.
+Algorithm 3, i.e., “Improved Ader”, is simply the result of applying the $\mathtt{FTS}$ meta-algorithm to Ader.
+We may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{IA})}({\mathbf{u}})[1,T]$ $=O\left(B_{0}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{3/4}\right),$ (1)
+where ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$.
+This matches the SOTA result of (?) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback.
+Moreover, by applying Theorem 8 to Improved Ader, we recover another result of (?)
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{IA})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{1/2}\right),$
+Online Gradient Descent with Separation Oracles: Similarly, the $\mathtt{SO\textrm{-}OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback obtains a weakly adaptive regret bound of $O(M_{1}T^{1/2})$ (Theorem 14 in (?)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
+$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(B_{0}T^{3/4}\right).$ (2)
+This matches the result of Theorem 15 in (?) and generalizes it to allow for stochastic feedback. Moreover, using Theorem 8, we see that
+$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(M_{1}T^{1/2}\right),$ (3)
+which is a novel result for adaptive regret of convex functions with deterministic zeroth order feedback.
 Remark 4.
-While $\operatorname{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\operatorname{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm.
-However, we note that even when ${\mathcal{A}}$ is semi-bandit, $\operatorname{FOTZO}({\mathcal{A}})$ is not a bandit algorithm.
-Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action.
-This is because $\operatorname{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm ${\mathcal{A}}$, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
-Remark 4.
-We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough.
-However, we have included the $\operatorname{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
-Recall that OGD algorithm Zinkevich, (2003) is a deterministic algorithm designed for deterministic semi-bandit feedback.
-Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result.
-This generalizes the result of Flaxman et al., (2005) to allow the feedback of the value oracle to be stochastic.
-Corollary 3.
-We have
-${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(DB_{0}T^{3/4}\right).$
-As mentioned in Remark 2, the Improved Ader algorithm of Zhang et al., 2018a is a deterministic algorithm designed for deterministic semi-bandit feedback.
-Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result.
-This matches the SOTA result of Zhao et al., (2021) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback.
-Corollary 4.
-Let IA denote the algorithm mentioned above.
-Also let ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$.
-Then we have
-${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{IA})}({\mathbf{u}})[1,T]$ $=O\left(DB_{0}\left(1+\frac{P_{T}({\mathbf{u}})}{D}\right)^{1/2}T^{3/4}\right).$
-Similarly, the SO-OGD algorithm Garber and Kretzu, (2022) is a deterministic algorithm designed for deterministic semi-bandit feedback with obtains a weakly adaptive regret bound of $O(DM_{1}T^{1/2})$ (Theorem 14 in Garber and Kretzu, (2022)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result.
-This matches the result of Theorem 15 in Garber and Kretzu, (2022) and generalizes it to allow for stochastic feedback.
-Corollary 5.
-Let SO-OGD be the algorithm mentioned above. Then we have
-$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{SO-OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(DB_{0}T^{3/4}\right).$
-10 Acknowledgements
+A careful review of Algorithms 6 and 7 in (?), together with their Lemmas 12 and 13 and Theorems 14 and 15 reveals that they added an extra layer of complexity to their base algorithm in order to be able to move from semi-bandit to bandit feedback.
+Specifically, if we instead use our Theorem 7, we can just drop $\delta^{\prime}$ from Algorithms 6 and 7 and Lemmas 12 and 13 (by setting it to zero) to obtain simpler and clearer algorithms, statements and proofs.
+Then their Theorem 15 will be an immediate corollary of applying our Theorem 7 to their previous results and the 3 pages of proof in their Appendix I will not be needed.
+12 Conclusions
+This paper presents a comprehensive framework for the analysis of meta-algorithms in online learning, offering a simplified approach to understanding and extending existing results in the field.
+By demonstrating the equivalence between online linear and convex optimization algorithms, particularly in the context of semi-bandit feedback, we provide a versatile tool for addressing a wide range of optimization problems.
+Moreover, our findings pave the way for the development of novel meta-algorithms, enabling the seamless conversion of algorithms across different settings.
+In addition to recovering many results in the literature with simplified proofs, we show that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$ with deterministic zeroth order feedback.
+This work opens various research opportunities to delve into the interplay between online linear and convex optimization algorithms.
+While it streamlines results across numerous scenarios, we anticipate that this approach will yield novel insights into some unresolved problems moving forward.
+13 Acknowledgments
 The authors want to acknowledge the useful feedback on the paper and its ideas provided by Yuanyu Wan.
-Appendix A Proof of Theorem 1
+A Proof of Theorem 1
 Proof.
@@
 ${\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}^{\prime}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}^{\prime}}={\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}^{{\mathcal{A}}}.\qed$
-Appendix B Proof of Theorem 2
+B Proof of Theorem 2
 Proof.
@@
 High probability regret bound:
-Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 1, we see that
+Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 4, we see that
 ${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}=\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\leq\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right)={\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}},$
@@
 ∎
-Appendix C Proof of Theorem 3
+C Proof of Theorem 3
 Proof.
@@
 which completes the proof. ∎
-Appendix D Proof of Theorem 4
+D Proof of Theorem 4
 Proof.
@@
 The proof of the claim for high probability regret bounds follows from Equation 5 and the same argument used in the last part of the proof of Theorem 2. ∎
-Appendix E Proof of Theorem 5
+E Proof of Theorem 5
 Proof.
@@
 We also define ${\mathcal{B}}^{\prime}_{\theta}:=({\mathcal{B}}^{\prime}_{\theta,1},\cdots,{\mathcal{B}}^{\prime}_{\theta,T})$.
-Note that a specific choice of $\theta$ is necessary to make sure that the function returned by ${\mathcal{B}}^{\prime}_{\theta,t}$ is a deterministic function of ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$ and not a random variable and therefore ${\mathcal{B}}^{\prime}_{\theta}$ belongs to $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])$.
+Note that a specific choice of $\theta$ is necessary to make sure that the function returned by ${\mathcal{B}}^{\prime}_{\theta,t}$ is a deterministic function of ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$ and not a random variable and therefore ${\mathcal{B}}^{\prime}_{\theta}$ belongs to $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])$.
 Since the algorithm uses (semi-)bandit feedback, given a specific value of $\theta$, the sequence of random vectors $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ chosen by ${\mathcal{A}}$ is identical between the game with ${\mathcal{B}}_{\theta}$ and ${\mathcal{B}}^{\prime}_{\theta}$.
@@
 Hence we have
-${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
+${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
 High probability regret bound:
@@
 ${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}_{\omega}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathbb{E}}_{\Omega^{\mathcal{Q}}}\left[{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\ \right]$
-$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}_{\omega}}$
+$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}_{\omega}}$
 Therefore we may use the same argument in the proof of Theorem 2 to see that
-¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
-Appendix F Proof of Theorems 6 and 7
-First we prove the following lemma which is a simple generalization of Observation 3 in Flaxman et al., (2005) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
+¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
+F Proof of Theorems 6 and 7
+First we prove the following lemma which is a simple generalization of Observation 3 in (?) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
 Lemma 1.
@@
 Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$.
-According to Lemma 7 in Pedramfar et al., (2023), we have ${\mathbf{z}}\in{\mathcal{K}}$.
+According to Lemma 7 in (?), we have ${\mathbf{z}}\in{\mathcal{K}}$.
 We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$.
@@
 where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$.
-Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in Pedramfar et al., (2023), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.
-[^3] When using a spherical estimator, it was shown in Flaxman et al., (2005) that $\hat{f}$ is differentiable even when $f$ is not.
-When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in Pedramfar et al., (2023).
+Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.
+[^3] When using a spherical estimator, it was shown in (?) that $\hat{f}$ is differentiable even when $f$ is not.
+When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in (?).
 However, their proof is based on the proof for the spherical case and therefore the results carry forward to show that $\hat{f}$ is differentiable.
 Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})$.
-Using Equation 3 and the definition of the Algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+Using Equation 6 and the definition of the Meta-algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
 It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the sequence of actions $({\mathbf{x}}_{1}+\delta{\mathbf{v}}_{1},\cdots,{\mathbf{x}}_{T}+\delta{\mathbf{v}}_{T})$ in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
 $\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}+\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T.$
-To prove the result in the Lipschitz case, we note that, according to Lemma 3 in Pedramfar et al., (2023), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$.
+To prove the result in the Lipschitz case, we note that, according to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$.
 By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
@@
 $\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
-$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in Pedramfar et al., (2023))
+$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
 $=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
@@
 As before, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
-It follows from the description of Algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+It follows from the description of Meta-algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
 Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$.
@@
 $={\mathbb{E}}\left[\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})\right)+\left(\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right].$ (10)
-To obtain the same bound as before, we note that in the convex case, instead of Inequality 5, we have
+To obtain the same bound as before, we note that in the convex case, instead of Inequality 8, we have
 $|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\frac{2M_{0}\delta}{\alpha}<\frac{4M_{0}\delta}{\alpha},$ (11)
-and, in the Lipschitz case, instead of Inequality F, we have
+and, in the Lipschitz case, instead of Inequality 9, we have
 $|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\delta M_{1}<2\delta M_{1}.$
 The rest of the proof follows verbatim. ∎
+G Proof of Theorem 8
+This proof is similar to the proofs of Theorems 6 and 7.
+Proof.
+Note that any realized adversary ${\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})$ may be represented as a sequence of functions $(f_{1},\cdots,f_{T})$.
+For such realized adversary ${\mathcal{B}}$, we define $\hat{B}$ to be the realized adversary corresponding to $(\hat{f}_{1},\cdots,\hat{f}_{T})$ with the stochastic gradient oracles
+$\hat{{\mathcal{Q}}}_{t}({\mathbf{x}}):=\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}},$ (12)
+where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$.
+Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Lemma 5 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.
+Moreover, we have
+$\|\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}}\|\leq\frac{k}{2\delta}M_{1}\|({\mathbf{x}}+\delta{\mathbf{v}})-({\mathbf{x}}-\delta{\mathbf{v}})\|\leq kM_{1}.$
+Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},kM_{1})$.
+Using Equation 12 and the definition of the Meta-algorithm 4, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$.
+We have
+${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $={\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]-{\mathbb{E}}\left[\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})\right]$
+$={\mathbb{E}}\left[\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})\right)+\left(\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right].$ (13)
+According to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. On the other hand, we have
+$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
+$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
+$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\delta}{r}\right){\mathbf{u}}_{t}+\frac{\delta}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
+$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\delta}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
+$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\delta M_{1}D}{r}\right)$ (Lipschitz)
+$=\left(1+\frac{2D}{r}\right)\delta M_{1}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
+Therefore, using Equation 13, we see that
+${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $\leq\delta M_{1}T+\left(1+\frac{2D}{r}\right)\delta M_{1}T=\left(2+\frac{2D}{r}\right)\delta M_{1}T.$
+Therefore, we have
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}$
+$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(2+\frac{2D}{r}\right)\delta M_{1}T$
+$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(2+\frac{2D}{r}\right)\delta M_{1}T$
+$\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}+\left(2+\frac{2D}{r}\right)\delta M_{1}T.\qed$

### v4

@@
 A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization
-Abstract
+Abstract.
 In this paper, we analyze the problem of online convex optimization in different settings, including different feedback types (full-information/semi-bandit/bandit/etc) in either stochastic or non-stochastic setting and different notions of regret (static adversarial regret/dynamic regret/adaptive regret).
 This is done through a framework which allows us to systematically propose and analyze meta-algorithms for the various settings described above.
-We show that any algorithm for online linear optimization with fully adaptive adversaries is an algorithm for online convex optimization.
+We show that any algorithm for online linear optimization with deterministic gradient feedback against fully adaptive adversaries is an algorithm for online convex optimization.
 We also show that any such algorithm that requires full-information feedback may be transformed to an algorithm with semi-bandit feedback with comparable regret bound.
@@
 Our framework allows us to analyze online optimization in various settings, recovers several results in the literature with a simplified proof technique, and provides new results.
-1 Introduction
-1.1 Overview
-Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information (?, ?).
+Key words and phrases:
+Online Convex Optimization; Meta-algorithms; Regret Minimization; Bandit Feedback; Gradient Estimation
+1. Introduction
+Online optimization problems represent a class of problems where the decision-maker, referred to as an agent, aims to make sequential decisions in the face of an adversary with incomplete information shalev-shwartz12_onlin_learn_onlin_convex_optim; hazan2016introduction.
 This setting mirrors a repeated game, where the agent and the adversary engage in a strategic interplay over a finite time horizon, commonly denoted as $T$.
@@
 Further, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries.
-Finally, we introduce meta algorithms, based on the variant of classical result of (?, ?), that convert algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit or zeroth-order feedback.
+Finally, we introduce meta algorithms, based on the variant of classical result of flaxman2005online; shamir17_optim_algor_bandit_zero_order, that convert algorithms for semi-bandit feedback, such as the ones mentioned above, to algorithms for bandit or zeroth-order feedback.
 In addition to recovering many results in the literature with simplified proofs, we give new static regret guarantees for online strongly convex optimization and adaptive regret guarantees for convex optimization with deterministic zeroth order feedback.
-Figure 1: Summary of the main results
+Figure 1. Summary of the main results
 Each arrow describes a procedure or meta-algorithm to convert an algorithm from one setting into another.
@@
 Finally, the last D or S, if exists, is used to denotes deterministic or stochastic algorithm.
-1.2 Contributions
 The key results of this paper are encapsulated in Figure 1, where each arrow represents a procedure or meta-algorithm that transforms an algorithm from one setting to another.
 The primary contribution is three-fold:
-  • (I)
-    We provide a framework that simplifies analysis of meta-algorithms in online learning.
-  • (II)
-    Using this framework, we formulate several meta-algorithms that facilitate the conversion of algorithms across various settings.
-  • (III)
-    In particular, we convert several insights in the literature into precise theorems, for example the idea of using 1-point gradient estimator in online optimization is formalized in Meta-algorithms 2 and 3 and Theorems 6 and 7.
-Similarly, the insight that “knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere” (See (?)) is formalized in Meta-algorithm 4 and Theorem 8.
+(I) We provide a framework that simplifies analysis of meta-algorithms in online learning.
+(II) Using this framework, we formulate several meta-algorithms that facilitate the conversion of algorithms across various settings.
+(III) In particular, we convert several insights in the literature into precise theorems, for example the idea of using 1-point gradient estimator in online optimization is formalized in Meta-algorithms 2 and 3 and Theorems 10 and 11.
+Similarly, the insight that “knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere” (See agarwal2010optimal) is formalized in Meta-algorithm 4 and Theorem 14.
 This framework not only simplifies proof analysis for numerous existing problems in the field (see Section 11) but also provides a tool for easier extension of multiple results in the literature.
@@
 (i) The first key result in this paper is Theorem 5.
-This result shows that any algorithm for online linear optimization with stochastic subgradient feedback results in a comparable regret with if the function class is replaced by a class of convex functions.
+This result shows that any algorithm for online linear optimization with deterministic subgradient feedback results in a comparable regret when the function class is replaced by a class of convex functions.
 This is the key result that allows us to go from online linear optimization to convex optimization in Figure 1 with semi-bandit feedback.
-We note that this result recovers the result for online convex game in (?), where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
+We note that this result recovers the result for online convex game in abernethy2008optimal, where the authors analyzed online convex optimization with deterministic gradient feedback and fully adaptive adversary and showed that the min-max regret bounds of online convex optimization is equal to that of online linear optimization (or online quadratic optimization is all convex functions are $\mu$-strongly convex).
 While the above result is the key for going from online linear to online convex optimization, the main issue for its applicability is that it requires semi-bandit feedback.
@@
 1.
-This result gives Theorem 6.5 in (?) as a special case.
+This result gives Theorem 6.5 in hazan2016introduction as a special case.
 Further, this allows us to have a result that generalizes both adversarial regret and stochastic regret, as they are commonly defined in the literature.
@@
 However, in many practical setups, we only have access to zeroth order or bandit feedback.
-In the next key result, given in Theorems 6-7, we use the smoothing trick (?, ?, ?), in order to estimate the gradient from the value oracle.
+In the next key result, given in Theorems 10-11, we use the smoothing trick nemirovsky83_probl_compl_method_effic_optim; flaxman2005online; pedramfar23_unified_approac_maxim_contin_dr_funct, in order to estimate the gradient from the value oracle.
 Using this estimated gradient, the results for zeroth order and bandit feedback are derived from that for first order and semi-bandit feedback, respectively.
 These results apply to any optimization problem, not necessarily convex.
-This approach recovers the results in (?) with the base algorithm of Online Gradient Descent ($\mathtt{OGD}$).
-In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback (?, ?, ?) with a simplified proof.
+This approach recovers the results in flaxman2005online with the base algorithm of Online Gradient Descent ($\mathtt{OGD}$).
+In particular, this approach also matches the SOTA dynamic and adaptive regret bounds for online convex optimization with bandit feedback zhao21_bandit_convex_optim_non_envir; garber22_new_projec_algor_onlin_convex; agarwal2010optimal with a simplified proof.
 (v) In the above results, we assumed that we only have access to an unbiased estimate of the zeroth order (or bandit) feedback.
-In (?), the setting of multi-point bandit feedback is introduced.
+In agarwal2010optimal, the setting of multi-point bandit feedback is introduced.
 Even though their setting and notion of regret is different from our work, one insight is relevant.
 Their analysis suggests that, in some cases, if we have access to exact values of the functions, knowing the value of each loss function at two points is almost as useful as knowing the value of each function everywhere.
-We formalize this insight by designing a meta-algorithm based on the two-point gradient estimator (?, ?) that allows us to convert algorithms for stochastic first order feedback into algorithms for deterministic zeroth order feedback with the same order of regret bound (Theorem 8).
+We formalize this insight by designing a meta-algorithm based on the two-point gradient estimator agarwal2010optimal; shamir17_optim_algor_bandit_zero_order that allows us to convert algorithms for stochastic first order feedback into algorithms for deterministic zeroth order feedback with the same order of regret bound (Theorem 14).
 Note that this result holds for convex and non-convex optimization as it only relies on the function class being Lipschitz.
-This result is shown to recover the result for dynamic regret for Improved Ader (?), while achieving two new results for adaptive regret for convex function of $O(\sqrt{T})$ and static regret for strongly convex function of $O(\log T)$.
+This result is shown to recover the result for dynamic regret for Improved Ader zhao21_bandit_convex_optim_non_envir, while achieving two new results for adaptive regret for convex function of $O(\sqrt{T})$ and static regret for strongly convex function of $O(\log T)$.
 Our work sheds light on the relation between different problems, algorithms and techniques used in online optimization and provides general meta-algorithms that allow conversion between different problem settings.
@@
 This broader applicability not only simplifies the analysis of numerous existing results within the field but also enables the derivation of multiple new results.
-For instance, results in (?, ?, ?, ?, ?, ?) as mentioned above can be demonstrated as corollaries of these more generalized results.
-Further, we provide new results, which shows that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$.
-1.3 Related works
-The term “online convex optimization” was first defined in (?), where Online Gradient Descent ($\mathtt{OGD}$) was proposed, while this setting was first introduced in (?).
-The Follow-The-Perturbed-Leader ($\mathtt{FTPL}$) for online linear optimization was introduced by (?), inspired by Follow-The-Leader algorithm originally suggested by (?).
-Follow-The-Regularized-Leader ($\mathtt{FTRL}$) in the context of online convex optimization was coined in (?, ?).
-The equivalence of $\mathtt{FTRL}$ and Online Mirror Descent ($\mathtt{OMD}$) was observed by (?).
-Non-stationary measures of regret have been studied in  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and many other papers.
-We refer the reader to surveys (?) and (?) for more details.
-2 Background and Notations
-For a set ${\mathcal{D}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{D}})$ to be the set of $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}$ for all ${\mathbf{x}},{\mathbf{y}}$ in ${\mathcal{D}}$ and $\alpha\in{\mathbb{R}}$.
-The relative interior of ${\mathcal{D}}$ is defined as
-$\operatorname{relint}({\mathcal{D}}):=\{{\mathbf{x}}\in{\mathcal{D}}\mid\exists r>0,{\mathbb{B}}_{r}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{D}})\subseteq{\mathcal{D}}\}.$
-We use ${\mathbb{B}}_{r}({\mathcal{D}})$ to denote the union of balls of radius $r$ centered at ${\mathcal{D}}$.
+For instance, results in abernethy2008optimal; hazan2016introduction; flaxman2005online; zhao21_bandit_convex_optim_non_envir; agarwal2010optimal; shamir17_optim_algor_bandit_zero_order as mentioned above can be demonstrated as corollaries of these more generalized results.
+Further, we provide new results for deterministic zeroth-order feedback case, showing that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$.
+Related works: The term “online convex optimization” was first defined in zinkevich03_onlin, where Online Gradient Descent ($\mathtt{OGD}$) was proposed, while this setting was first introduced in gordon99_regret.
+The Follow-The-Perturbed-Leader ($\mathtt{FTPL}$) for online linear optimization was introduced by kalai05_effic, inspired by Follow-The-Leader algorithm originally suggested by hannan57.
+Follow-The-Regularized-Leader ($\mathtt{FTRL}$) in the context of online convex optimization was coined in shalev-shwartz07; shalev-shwartz07_onlin.
+The equivalence of $\mathtt{FTRL}$ and Online Mirror Descent ($\mathtt{OMD}$) was observed by hazan10_extrac.
+Non-stationary measures of regret have been studied in zinkevich03_onlin; hazan09_effic; besbes15_non_station_stoch_optim; daniely15_stron_adapt_onlin_learn; zhang18_dynam_regret_stron_adapt_method; zhang18_adapt_onlin_learn_dynam_envir; zhao20_dynam_regret_convex_smoot_funct; zhao21_bandit_convex_optim_non_envir; lu23_projec_adapt_regret_member_oracl; wang24_non_projec_free_onlin_learn; garber22_new_projec_algor_onlin_convex and many other papers.
+We refer the reader to surveys shalev-shwartz12_onlin_learn_onlin_convex_optim and hazan2016introduction for more details.
+2. Background and Notations
+For a set ${\mathcal{S}}\subseteq{\mathbb{R}}^{d}$, we define its affine hull $\operatorname{aff}({\mathcal{S}})$ to be the set of all points of the form $\sum_{i=1}^{m}\alpha_{i}{\mathbf{x}}_{i}$ for integer $m\geq 1$, points $({\mathbf{x}}_{i})_{i=1}^{m}$ in $\mathcal{S}$}, and real numbers $(\alpha_{i})_{i=1}^{m}$.
+The relative interior of $\mathcal{S}$} is defined as
+$\operatorname{relint}({\mathcal{S}}):=\{{\mathbf{x}}\in{\mathcal{S}}\mid\exists r>0,{\mathbb{B}}_{r}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{S}})\subseteq{\mathcal{S}}\}.$
+We use ${\mathbb{B}}_{r}({\mathcal{S}})$ to denote the union of balls of radius $r$ centered at $\mathcal{S}$} and use $\operatorname{diam}({\mathcal{S}}):=\sup_{{\mathbf{x}},{\mathbf{y}}\in{\mathcal{S}}}\|{\mathbf{x}}-{\mathbf{y}}\|$ to denote the diameter of $\mathcal{S}$}.
 A function class is simply a set of real-valued functions.
-Given a set ${\mathcal{D}}$, a function class over ${\mathcal{D}}$ is simply a subset of all real-valued functions on ${\mathcal{D}}$.
-Given a real number $\mu\geq 0$ and a set ${\mathcal{D}}$, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{D}}$.
+Given a set $\mathcal{S}$}, a function class over $\mathcal{S}$} is simply a subset of all real-valued functions on $\mathcal{S}$}.
+Given a real number $\mu\geq 0$ and a set $\mathcal{S}$}, we use ${\mathbf{Q}}_{\mu}$ to denote the class of functions of the form $q_{{\mathbf{x}},{\mathbf{o}}}({\mathbf{y}})=\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, where ${\mathbf{o}}\in{\mathbb{R}}^{d}$ and ${\mathbf{x}}\in{\mathcal{S}}$.
 A set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is called a convex set if for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{K}}$ and $\alpha\in[0,1]$, we have $\alpha{\mathbf{x}}+(1-\alpha){\mathbf{y}}\in{\mathcal{K}}$.
@@
 A vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is called a subgradient of $f$ at $\mathbf{x}$} if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle$, for all ${\mathbf{y}}\in{\mathcal{K}}$.
-Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex if and only if it has a subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$.
 More generally, given $\mu\geq 0$, we say a vector ${\mathbf{o}}\in{\mathbb{R}}^{d}$ is a $\mu$-subgradient of $f$ at $\mathbf{x}$} if $f({\mathbf{y}})-f({\mathbf{x}})\geq\langle{\mathbf{o}},{\mathbf{y}}-{\mathbf{x}}\rangle+\frac{\mu}{2}\|{\mathbf{y}}-{\mathbf{x}}\|^{2}$, for all ${\mathbf{y}}\in{\mathcal{K}}$.
-Given a convex set ${\mathcal{K}}$, a function $f:{\mathcal{K}}\to{\mathbb{R}}$ is $\mu$-strongly convex if and only if it has a $\mu$-subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$.
+Given a convex set $\mathcal{K}$}, a Lipschitz continuous function $f:{\mathcal{K}}\to{\mathbb{R}}$ is convex (resp.
+$\mu$-strongly convex) if and only if it has a ($\mu$-)subgradient at all points ${\mathbf{x}}\in{\mathcal{K}}$.
 We use $\nabla^{*}f$ to denote the set of $\mu$-subgradients of $f$.
-3 Problem Setup
+3. Problem Setup
 Online optimization problems can be formalized as a repeated game between an agent and an adversary.
@@
 Formally, a query oracle for a function $f$ is a map $\mathcal{Q}$} defined on $\mathcal{K}$} such that for each ${\mathbf{x}}\in{\mathcal{K}}$, the ${\mathcal{Q}}({\mathbf{x}})$ is a random variable taking value in the observation space $\mathcal{O}$}.
-The query oracle is called a stochastic value oracle or stochastic zeroth order oracle if ${\mathcal{O}}={\mathbb{R}}$ and $f({\mathbf{x}})={\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$.
+The query oracle is called a stochastic value oracle or stochastic zeroth order oracle if $\mathcal{O}$$\mathbb{R}$} and $f({\mathbf{x}})={\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$.
 Similarly, it is called a stochastic (sub)gradient oracle or stochastic first order oracle if ${\mathcal{O}}={\mathbb{R}}^{d}$ and ${\mathbb{E}}[{\mathcal{Q}}({\mathbf{x}})]$ is a (sub)gradient of $f$ at $\mathbf{x}$}.
@@
 Similarly, it takes bandit feedback if the oracle is zeroth-order and the agent query function is trivial  [^1] This is a slight generalization of the common use of the term bandit feedback.
-Usually, bandit feedback refers to the case where the oracle is a deterministic zeroth-order oracle and the agent query function is trivial..
+Usually, bandit feedback refers to the case where the oracle is a deterministic zeroth-order oracle and the agent query function is trivial, while our definition allows for stochastic oracles..
 If the agent query function is non-trivial, then we say the agent requires full-information feedback.
@@
 In this case, a realized adversary may be simply represented by a sequence of functions $(f_{1},\cdots,f_{T})\in{\mathbf{F}}^{T}$ and a sequence of query oracles $({\mathcal{Q}}_{1},\cdots,{\mathcal{Q}}_{T})$ for these functions.
-We say an adversary is a weakly adaptive adversary if each function ${\mathcal{B}}_{t}$ described above does not depend on ${\mathbf{x}}_{t}$ and therefore may be represented as a map defined on ${\mathcal{K}}^{t-1}$.
 In this work we also consider adversaries that are fully adaptive, i.e., adversaries with no restriction.
+[^3] Another form of adversary considered in literature is a weakly adaptive adversary where each function ${\mathcal{B}}_{t}$ described above does not depend on ${\mathbf{x}}_{t}$ and therefore may be represented as a map defined on ${\mathcal{K}}^{t-1}$.
 Clearly any oblivious adversary is a weakly adaptive adversary and any weakly adaptive adversary is a fully adaptive adversary.
@@
 In order to handle different notions of regret with the same approach, for an agent $\mathcal{A}$}, adversary $\operatorname{Adv}$, compact set ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ and $1\leq a\leq b\leq T$, we define regret as
-${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ $:=\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\min_{{\mathbf{u}}=({\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{T})\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right],$
-where the expectation in the definition of the regret is over the randomness of the algorithm and the query oracles.
+${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ $:=\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]$
+where ${\mathbf{u}}=({\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{T})$ and the expectation in the definition of the regret is over the randomness of the algorithm and the query oracles.
 We use the notation ${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]:={\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ when $\operatorname{Adv}=\{{\mathcal{B}}\}$ is a singleton.
 Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$.
-When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret (?, ?).
-If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret (?).
-Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ (?).
+When $a=1$, $b=T$ and $\mathcal{U}$} contains only a single element then it is referred to as the dynamic regret zinkevich03_onlin; zhang18_adapt_onlin_learn_dynam_envir.
+This notion of regret allows us to provide an upper bound on regret that may depend on the comparator.
+In Section 11, we will discuss algorithms that provide guarantees depending on the path-length, defined with $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$.
+Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ hazan09_effic.
+A crucial point in this definition is that the comparison is with respect to a (potentially) different optimum for any interval.
 We drop $a$, $b$ and $\mathcal{U}$} when the statement is independent of their value or their value is clear from the context.
@@
 We say $h:[0,1]\to[0,\infty)$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$ if for each ${\mathcal{B}}\in\operatorname{Adv}$ and $\delta\in[0,1]$, we have
-(${\mathbb{P}}\left(\{\omega\in\Omega\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
+${\mathbb{P}}\left(\{\omega\in\Omega\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
 Note that the infimum of any family of high probability regret bounds is a high probability regret bound.
-Hence we use ¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}}:[0,1]\to[0,\infty)$ to denote the smallest high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$.
-In other words, ¯$\overline{{\mathcal{R}}}_{{\mathcal{B}}}^{{\mathcal{A}}}$ is the quantile function for the random variable $\omega\mapsto{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}$.
-4 Re-statement of Previous Result: Oblivious to fully adaptive adversary
+Hence we use $\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}}:[0,1]\to[0,\infty)$ to denote the smallest high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$.
+In other words, $\overline{{\mathcal{R}}}_{{\mathcal{B}}}^{{\mathcal{A}}}$ is the quantile function for the random variable $\omega\mapsto{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}$.
+4. Re-statement of Previous Result: Oblivious to fully adaptive adversary
 The following theorem states a result that is well-known in the literature.
-For example, (?) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary.
+For example, hazan2016introduction only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary.
 Here we explicitly mention it as a theorem for completion.
@@
 The $\mathtt{OGD}$, $\mathtt{FTRL}$, and $\mathtt{OMD}$ algorithms are deterministic and therefore have same regret bound in fully-adaptive setting as the oblivious setting.
-5 Linear to convex with fully adaptive adversary
+5. Linear to convex with fully adaptive adversary
 Here we show that any semi-bandit feedback online linear optimization algorithm for fully adaptive adversary is also an online convex optimization algorithm. We start with a definition.
@@
 See Appendix B for proof.
-Corollary 2 (Theorem 14 in (?)).
-The min-max regret bound of the problem of online linear optimization with deterministic feedback and online convex optimization with deterministic feedback are equal.
+Corollary 0 (Theorem 14 in abernethy2008optimal).
+The min-max regret bounds of online linear optimization with deterministic feedback and online convex optimization with deterministic feedback are equal.
 To see why this is true, note that any algorithm has the same performance in both settings.
 So it follows that an optimal algorithm for one setting is optimal for the other and therefore the min-max regret bounds are equal.
-6 Full information feedback to semi-bandit feedback
-Any algorithm designed for semi-bandit setting may be trivially applied in the first-order full-information feedback setting.
-In the following theorem, we show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough.
-Specifically, we show that an algorithm that requires full information feedback could be converted to an algorithm that only requires semi-bandit feedback with the same regret bounds.
-The meta-algorithm that does this conversion is described in Meta-algorithm 1.
-Next we show that $\mathtt{FTS}({\mathcal{A}})$ always performs at least as good as ${\mathcal{A}}$ when the oracle is deterministic.
+6. Full information feedback to semi-bandit feedback
 Input : horizon $T$, algorithm ${\mathcal{A}}=({\mathcal{A}}^{\text{action}},{\mathcal{A}}^{\text{query}})$, strong convexity coefficient $\mu\geq 0$
@@
      Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
-             Return $\nabla q_{t}({\mathbf{y}}_{t,i})={\mathbf{o}}_{t}+\mu\langle{\mathbf{y}}_{t,i},{\mathbf{y}}_{t,i}-{\mathbf{x}}_{t}\rangle$ as the output of the query oracle to ${\mathcal{A}}$
+     Return $\nabla q_{t}({\mathbf{y}}_{t,i})={\mathbf{o}}_{t}+\mu({\mathbf{y}}_{t,i}-{\mathbf{x}}_{t})$ as the output of the query oracle to $\mathcal{A}$}
   end for
@@
 Meta-algorithm 1 Full information to semi-bandit - $\mathtt{FTS}({\mathcal{A}})$
+Any algorithm designed for semi-bandit setting may be trivially applied in the first-order full-information feedback setting.
+In the following theorem, we show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough.
+Specifically, we show that an algorithm that requires full information feedback could be converted to an algorithm that only requires semi-bandit feedback with the same regret bounds.
+The meta-algorithm that does this conversion is described in Meta-algorithm 1.
+Next we show that $\mathtt{FTS}({\mathcal{A}})$ always performs at least as good as $\mathcal{A}$} when the oracle is deterministic.
 Theorem 7.
@@
 If $\mathbf{F}$} is a $\mu$-strongly convex function class over $\mathcal{K}$} , then we have
-${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
+${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}},$
 where ${\mathcal{A}}^{\prime}=\mathtt{FTS}({\mathcal{A}})$ is Meta-algorithm 1. In particular, if $\mathbf{F}$} is closed under $\mu$-quadratization, then
-${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
+${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}.$
 See Appendix C for proof. Note that when the base algorithm $\mathcal{A}$} is semi-bandit, we have $\mathtt{FTS}({\mathcal{A}})={\mathcal{A}}$ and the above theorem becomes trivial.
-7 Lipschitz to non-Lipschitz
-It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in (?) and Lemma 1).
-Here we use a shrinking method described in (?).
+7. Lipschitz to non-Lipschitz
+It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in flaxman2005online and Lemma 16).
+Here we use a shrinking method described in pedramfar23_unified_approac_maxim_contin_dr_funct.
 We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}({\mathbf{c}})\cap\operatorname{aff}({\mathcal{K}})\subseteq{\mathcal{K}}$.
@@
 Then, for any adversary $\operatorname{Adv}$, we have
-${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}.$
+${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})$ $\leq{\mathcal{R}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r},$
+$\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})$ $\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}.$
 See Appendix D for proof.
-8 Deterministic feedback to stochastic feedback
+8. Deterministic feedback to stochastic feedback
 In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries.
-In particular, there is no need for any variance reduction method, such as momentum (?, ?, ?), as long as we know that the algorithm is designed for a fully adaptive adversary.
+In particular, there is no need for any variance reduction method, such as momentum mokhtari20_stoch_condit_gradien_method; chen18_projec_free_onlin_optim_stoch_gradien; xie20_effic_projec_free_onlin_method, as long as we know that the algorithm is designed for a fully adaptive adversary.
 Theorem 9.
@@
 If $\mathbf{F}$} is an $M_{1}$-Lipschitz $\mu$-strongly convex function class over $\mathcal{K}$} and $B_{1}\geq M_{1}$, then we have
-${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
-See Appendix E for proof. Theorem 6.5 in (?) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
-9 First order feedback to zeroth order feedback
-When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient.
-The “smoothing trick” is a classical idea in convex optimization (See (?)) which was first used in online convex optimization in (?).
-This idea involves averaging through spherical sampling around a given point.
-Here we use a variant that was introduced in (?) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
+${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}$ $\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}},$
+$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}$ $\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}.$
+See Appendix E for proof. Theorem 6.5 in hazan2016introduction may be seen as a special case of the above theorem when $\mathcal{U}$} is a singleton and $\mathcal{A}$} is deterministic.
+9. First order feedback to zeroth order feedback
 Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm $\mathcal{A}$}
@@
      Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
-             Let $o_{t,i}$ be the output of the query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$
-             Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
+     Let $o_{t,i}$ be the output of the query oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$
+     Pass $\frac{k}{\delta}o_{t,i}{\mathbf{v}}_{t,i}$ as the oracle output to $\mathcal{A}$}
   end for
@@
 Meta-algorithm 2 First order to zeroth order - $\mathtt{FOTZO}({\mathcal{A}})$
+When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient.
+The “smoothing trick” is a classical idea in convex optimization (See nemirovsky83_probl_compl_method_effic_optim) which was first used in online convex optimization in flaxman2005online.
+This idea involves averaging through spherical sampling around a given point.
+Here we use a variant that was introduced in pedramfar23_unified_approac_maxim_contin_dr_funct which does not require extra assumptions on the convex set $\mathcal{K}$}.
+For a function $f:{\mathcal{K}}\to{\mathbb{R}}$ defined on a convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ and a smoothing parameter $0<\delta\leq\alpha$, its $\delta$-smoothed version $\hat{f}_{\delta}:\hat{{\mathcal{K}}}_{\alpha}\to{\mathbb{R}}$ is given as
+$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{K}})}[f({\mathbf{z}})]={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
 Input : Shrunk domain $\hat{{\mathcal{K}}}_{\alpha}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta\leq\alpha$, horizon $T$, algorithm $\mathcal{A}$}
@@
 Meta-algorithm 3 Semi-bandit to bandit - $\mathtt{STB}({\mathcal{A}})$
-For a function $f:{\mathcal{K}}\to{\mathbb{R}}$ defined on a convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ and a smoothing parameter $0<\delta\leq\alpha$, its $\delta$-smoothed version $\hat{f}_{\delta}:\hat{{\mathcal{K}}}_{\alpha}\to{\mathbb{R}}$ is given as
-$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{K}})}[f({\mathbf{z}})]$
-$={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
 where ${\mathcal{L}}_{0}=\operatorname{aff}({\mathcal{K}})-{\mathbf{x}}$, for any ${\mathbf{x}}\in{\mathcal{K}}$ is the linear space that is a translation of the affine hull of $\mathcal{K}$} and $\mathbf{v}$} is sampled uniformly at random from the $k=\operatorname{dim}({\mathcal{L}}_{0})$-dimensional ball ${\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}$.
@@
   • if $\mathbf{F}$} is $M_{1}$-Lipschitz (but not necessarily convex), then we have
-    ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T.$
+    ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})$ $\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T,$
+    where $D=\operatorname{diam}({\mathcal{K}})$.
 Theorem 11.
 Under the assumptions of Theorem 10, if we further assume that $\mathcal{A}$} is semi-bandit, then the same regret bounds hold with ${\mathcal{A}}^{\prime}=\mathtt{STB}({\mathcal{A}})$, where $\mathtt{STB}$ is described by Meta-algorithm 3.
-See Appendix F for proof.
-Remark 2.
-While $\mathtt{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\mathtt{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm.
-However, we note that even when ${\mathcal{A}}$ is semi-bandit, $\mathtt{FOTZO}({\mathcal{A}})$ is not a bandit algorithm.
-Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action.
-This is because $\mathtt{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm ${\mathcal{A}}$, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
-Remark 3.
-We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough.
-However, we have included the $\mathtt{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
-10 First order feedback to deterministic zeroth order feedback
-When we have access to a deterministic zeroth order feedback, we may use the two point gradient estimator (?, ?) to significantly improve the results of the previous section.
-Theorem 8.
-Let ${\mathbf{F}}$ be an $M_{1}$-Lipschitz function class over a convex set ${\mathcal{K}}$ and choose ${\mathbf{c}}$ and $r$ as described above and let $\delta<r$.
-Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\delta}{r}){\mathcal{U}}+\frac{\delta}{r}{\mathbf{c}}$.
-Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback.
-Then, if ${\mathcal{A}}^{\prime}=\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$ where $\mathtt{FOTZO\textrm{-}2P}$ is described by Meta-algorithm 2, we have
-${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D}{r}\right)\delta M_{1}T.$
-Note that here we may choose $\delta=T^{-1}$ to see that the order of regret of ${\mathcal{A}}^{\prime}$ remains the same as that of ${\mathcal{A}}$. See Appendix G for proof.
 Input : Shrunk domain $\hat{{\mathcal{K}}}_{\delta}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta<r$, horizon $T$, algorithm $\mathcal{A}$}
@@
 end for
-Meta-algorithm 4 First order to zeroth order with Two Point gradient estimator - $\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$
-11 Applications
+Meta-algorithm 4 First order to zeroth order with two-point gradient estimator - $\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$
+See Appendix F for proof.
+Remark 0.
+While $\mathtt{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\mathtt{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm.
+However, we note that even when $\mathcal{A}$} is semi-bandit, $\mathtt{FOTZO}({\mathcal{A}})$ is not a bandit algorithm.
+Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action.
+This is because $\mathtt{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm $\mathcal{A}$}, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
+Remark 0.
+We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough.
+However, we have included the $\mathtt{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
+10. First order feedback to deterministic zeroth order feedback
+When we have access to a deterministic zeroth order feedback, we may use the two point gradient estimator agarwal2010optimal; shamir17_optim_algor_bandit_zero_order to significantly improve the results of the previous section.
+Theorem 14.
+Let $\mathbf{F}$} be an $M_{1}$-Lipschitz function class over a convex set $\mathcal{K}$} and choose $\mathbf{c}$} and $r$ as described above and let $\delta<r$.
+Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\delta}{r}){\mathcal{U}}+\frac{\delta}{r}{\mathbf{c}}$.
+Assume $\mathcal{A}$} is an algorithm for online optimization with first order feedback.
+Then, if ${\mathcal{A}}^{\prime}=\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$ where $\mathtt{FOTZO\textrm{-}2P}$ is described by Meta-algorithm 4, we have
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(2+\frac{2D}{r}\right)\delta M_{1}T,$
+where $D=\operatorname{diam}({\mathcal{K}})$.
+Note that here we may choose $\delta=T^{-1}$ to see that the order of regret of $\mathcal{A}$$^{\prime}$ remains the same as that of $\mathcal{A}$}. See Appendix G for proof.
+11. Applications
 In this section, we discuss applications of meta-algorithms discussed above to some specific base algorithms.
 Note that all the results stated here are immediate applications of our framework to existing results and no extra proof is needed.
-Online Gradient Descent: Recall that $\mathtt{OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback that obtains a regret bound of $O\left(M_{1}T^{1/2}\right)$ over $M_{1}$-Lipschitz convex class ${\mathbf{F}}$ and $O\left(M_{1}^{2}\log(T)\right)$ over strongly convex $M_{1}$-Lipschitz convex class ${\mathbf{F}}^{\prime}$ (see (?)).
-Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$ and $\alpha=\delta=(T\log(T))^{-1/2}$, respectively, to see that
-${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{0}T^{3/4}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]=O\left(B_{0}^{2}T^{1/2}\log(T)^{1/2}\right).$
-The first result generalizes the result of (?) to allow the feedback of the value oracle to be stochastic.
-The second result provides a new optimal algorithm for bandit strongly convex optimization, matching the result of (?).
-However, (?) obtains a regret bound of the same order for the strongly convex case while assuming access to a self-concordant barrier for the domain and also assuming that the functions are smooth.
-We do not require either of these assumptions here.
-On the other hand, using Theorem 8 on $\mathtt{OGD}$, we see that:
-${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right),$
-where the first result is proven in (?) and the second one is novel.
-Ader and Improved Ader: In (?), Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback.
+Online Gradient Descent: Recall that $\mathtt{OGD}$ algorithm zinkevich03_onlin is a deterministic algorithm designed for deterministic semi-bandit feedback that obtains a regret bound of $O\left(M_{1}T^{1/2}\right)$ over $M_{1}$-Lipschitz convex class $\mathbf{F}$} and $O\left(M_{1}^{2}\log(T)\right)$ over strongly convex $M_{1}$-Lipschitz convex class $\mathbf{F}$$^{\prime}$ (see hazan2016introduction).
+[^4] Here we assume $\operatorname{diam}({\mathcal{K}})\leq 1$ for simplicity.
+General bounds follow similarly and may depend on $\operatorname{diam}({\mathcal{K}})$ and $r$, as described in Theorems 10 and 14.
+Since $\mathtt{OGD}$ is deterministic, using Theorem 2, we see that
+${\mathcal{R}}_{\operatorname{Adv}^{\text{f}}_{1}({\mathbf{F}})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),$
+$\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{f}}_{1}({\mathbf{F}}^{\prime})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right).$
+Hence we may apply Theorems 9 to see that, if we have access to stochastic gradient oracle bounded by $B_{1}\geq M_{1}$, then we have
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}},B_{1})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{1}T^{1/2}\right),$
+$\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}}^{\prime},B_{1})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{1}^{2}\log(T)\right).$
+For convex functions, we may apply Theorem 11, with $\alpha=\delta=T^{-1/4}$. Thus, if we have access to stochastic value oracle bounded by $B_{0}\geq M_{1}$, we have
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}},\frac{k}{\delta}B_{0})}^{\mathtt{OGD}}(\hat{{\mathcal{K}}}_{*}^{T})[1,T]+O\left(\delta M_{1}T\right)$
+$=O\left(\frac{k}{\delta}B_{0}T^{1/2}+\delta M_{1}T\right)=O\left(B_{0}T^{3/4}\right),$
+which generalizes the result of flaxman2005online to allow the feedback of the value oracle to be stochastic.
+Similarly, for strongly convex functions, we may apply Theorem 11, with $\alpha=\delta=T^{-1/3}\log(T)^{1/3}$ to see that
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}}^{\prime},\frac{k}{\delta}B_{0})}^{\mathtt{OGD}}(\hat{{\mathcal{K}}}_{*}^{T})[1,T]+O\left(\delta M_{1}T\right)$
+$=O\left(\frac{k^{2}}{\delta^{2}}B_{0}^{2}\log(T)+\delta M_{1}T\right)=O\left(B_{0}^{2}T^{2/3}\log(T)^{1/3}\right).$
+This provides a new algorithm for bandit strongly convex optimization, matching the result of agarwal2010optimal.
+However, if we further assume that the functions are smooth, then the regret bound could be improved to $\widetilde{O}(\sqrt{T})$ (See hazan14_bandit_convex_optim; ito20_optim_algor_bandit_convex_optim).
+On the other hand, using Theorem 14 on $\mathtt{OGD}$, we see that
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),$
+$\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right),$
+where the first result is proven in shamir17_optim_algor_bandit_zero_order and the second one is novel.
+Ader and Improved Ader: In zhang18_adapt_onlin_learn_dynam_envir, Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback.
 Algorithm 3, i.e., “Improved Ader”, is simply the result of applying the $\mathtt{FTS}$ meta-algorithm to Ader.
-We may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
+We may apply Theorems 9 and 11, with $\alpha=\delta=T^{-1/4}$, to see that
 ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{IA})}({\mathbf{u}})[1,T]$ $=O\left(B_{0}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{3/4}\right),$ (1)
 where ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$.
-This matches the SOTA result of (?) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback.
-Moreover, by applying Theorem 8 to Improved Ader, we recover another result of (?)
-${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{IA})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{1/2}\right),$
-Online Gradient Descent with Separation Oracles: Similarly, the $\mathtt{SO\textrm{-}OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback obtains a weakly adaptive regret bound of $O(M_{1}T^{1/2})$ (Theorem 14 in (?)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
-$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(B_{0}T^{3/4}\right).$ (2)
-This matches the result of Theorem 15 in (?) and generalizes it to allow for stochastic feedback. Moreover, using Theorem 8, we see that
-$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(M_{1}T^{1/2}\right),$ (3)
+This matches the SOTA result of zhao21_bandit_convex_optim_non_envir for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback.
+Moreover, by applying Theorem 14 to Improved Ader, we recover another result of zhao21_bandit_convex_optim_non_envir
+${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{IA})}({\mathbf{u}})[1,T]$ $=O\left(M_{1}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{1/2}\right).$
+Online Gradient Descent with Separation Oracles: Similarly, the $\mathtt{SO\textrm{-}OGD}$ algorithm garber22_new_projec_algor_onlin_convex is a deterministic algorithm designed for deterministic semi-bandit feedback obtains a adaptive regret bound of $O(M_{1}T^{1/2})$ (Theorem 14 in garber22_new_projec_algor_onlin_convex) Hence we may apply Theorems 9 and 11, with $\alpha=\delta=T^{-1/4}$, to see that
+$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]=O\left(B_{0}T^{3/4}\right).$
+This matches the result of Theorem 15 in garber22_new_projec_algor_onlin_convex and generalizes it to allow for stochastic feedback. Moreover, using Theorem 14, we see that
+$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]=O\left(M_{1}T^{1/2}\right),$
 which is a novel result for adaptive regret of convex functions with deterministic zeroth order feedback.
 Remark 0.
-A careful review of Algorithms 6 and 7 in (?), together with their Lemmas 12 and 13 and Theorems 14 and 15 reveals that they added an extra layer of complexity to their base algorithm in order to be able to move from semi-bandit to bandit feedback.
+A careful review of Algorithms 6 and 7 in garber22_new_projec_algor_onlin_convex, together with their Lemmas 12 and 13 and Theorems 14 and 15 reveals that they added an extra layer of complexity to their base algorithm in order to be able to move from semi-bandit to bandit feedback.
 Specifically, if we instead use our Theorem 7, we can just drop $\delta^{\prime}$ from Algorithms 6 and 7 and Lemmas 12 and 13 (by setting it to zero) to obtain simpler and clearer algorithms, statements and proofs.
 Then their Theorem 15 will be an immediate corollary of applying our Theorem 11 to their previous results and the 3 pages of proof in their Appendix I will not be needed.
-12 Conclusions
+12. Conclusions
 This paper presents a comprehensive framework for the analysis of meta-algorithms in online learning, offering a simplified approach to understanding and extending existing results in the field.
@@
 Moreover, our findings pave the way for the development of novel meta-algorithms, enabling the seamless conversion of algorithms across different settings.
-In addition to recovering many results in the literature with simplified proofs, we show that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$ with deterministic zeroth order feedback.
+In addition to recovering many results in the literature with simplified proofs, for convex function with deterministic zeroth order feedback, we obtain an adaptive regret of $O(\sqrt{T})$ (with $\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})$) and static regret of $O(\log T)$ for strongly convex case (with $\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})$).
 This work opens various research opportunities to delve into the interplay between online linear and convex optimization algorithms.
 While it streamlines results across numerous scenarios, we anticipate that this approach will yield novel insights into some unresolved problems moving forward.
-13 Acknowledgments
+13. Acknowledgments
 The authors want to acknowledge the useful feedback on the paper and its ideas provided by Yuanyu Wan.
-A Proof of Theorem 1
+Appendix A Proof of Theorem 2
 Proof.
@@
 ${\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{i}^{\text{f}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}^{\prime}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}{\mathcal{R}}^{\mathcal{A}}_{{\mathcal{B}}^{\prime}}={\mathcal{R}}_{\operatorname{Adv}_{i}^{\text{o}}({\mathbf{F}})}^{{\mathcal{A}}}.\qed$
-B Proof of Theorem 2
+Appendix B Proof of Theorem 5
 Proof.
@@
 High probability regret bound:
-Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 4, we see that
+Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 2, we see that
 ${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}=\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\leq\max_{{\mathbf{u}}\in{\mathcal{U}}}\left(\sum_{t=a}^{b}q_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}q_{t}({\mathbf{u}}_{t})\right)={\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}},$
 which immediately implies that
-(${\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
+${\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
 Hence $h$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}))$ as well.
@@
 ∎
-C Proof of Theorem 3
+Appendix C Proof of Theorem 7
 Proof.
-First we note that for any adversary $\operatorname{Adv}$ over ${\mathbf{Q}}_{\mu}$ with deterministic first order feedback, under the algorithms ${\mathcal{A}}^{\prime}$, we have $q_{t}=f_{t}$.
+First we note that for any adversary $\operatorname{Adv}$ over ${\mathbf{Q}}_{\mu}$ with deterministic first order feedback, under the algorithms $\mathcal{A}$$^{\prime}$, the functions $q_{t}$ and $f_{t}$ are equal up to an additive constant.
 Hence $\mathcal{A}$$^{\prime}$ and $\mathcal{A}$} return the same sequence of points and we have
@@
 and similarly
-¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}.$
+$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}.$
 As a special case, if $\mathbf{F}$} is closed under $\mu$-quadratization, then we may use the same argument in the proof of Theorem 5 to see that
@@
 and
-¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
+$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
 which completes the proof. ∎
-D Proof of Theorem 4
+Appendix D Proof of Theorem 8
 Proof.
@@
 Putting these together, we see that
-${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})-{\mathcal{R}}_{{\mathcal{B}}|_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})$ $\leq\frac{\alpha M_{0}T}{r}.$ (5)
+[…]
 Therefore, we have
@@
 The proof of the claim for high probability regret bounds follows from Equation 3 and the same argument used in the last part of the proof of Theorem 5. ∎
-E Proof of Theorem 5
+Appendix E Proof of Theorem 9
 Proof.
@@
 Therefore we may use the same argument in the proof of Theorem 5 to see that
-¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
-F Proof of Theorems 6 and 7
-First we prove the following lemma which is a simple generalization of Observation 3 in (?) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
+$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
+Appendix F Proof of Theorems 10 and 11
+First we prove the following lemma which is a simple generalization of Observation 3 in flaxman2005online to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
 Lemma 0.
@@
 Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$.
-According to Lemma 7 in (?), we have ${\mathbf{z}}\in{\mathcal{K}}$.
+According to Lemma 7 in pedramfar23_unified_approac_maxim_contin_dr_funct, we have ${\mathbf{z}}\in{\mathcal{K}}$.
 We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$.
@@
 $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}}):=\frac{k}{\delta}{\mathcal{Q}}_{t}({\mathbf{x}}+\delta{\mathbf{v}}){\mathbf{v}},$ (4)
-where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$.
-Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.
-[^3] When using a spherical estimator, it was shown in (?) that $\hat{f}$ is differentiable even when $f$ is not.
-When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in (?).
-However, their proof is based on the proof for the spherical case and therefore the results carry forward to show that $\hat{f}$ is differentiable.
-Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})$.
-Using Equation 6 and the definition of the Meta-algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
-It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the sequence of actions $({\mathbf{x}}_{1}+\delta{\mathbf{v}}_{1},\cdots,{\mathbf{x}}_{T}+\delta{\mathbf{v}}_{T})$ in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+[…]
 Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$.
@@
 On the other hand, we have
-$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
-$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 1)
-$=\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
-$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}f_{t}({\mathbf{c}})\right)$ (Convexity)
-$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}M_{0}\right)$
-$=\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\left(1-\frac{\alpha}{r}\right)\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
+$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{w}}}_{t})$
+$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{w}}}_{t})$ (Lemma 16)
+$=\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{w}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
+$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{w}}_{t})+\frac{\alpha}{r}f_{t}({\mathbf{c}})\right)$ (Convexity)
+$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{w}}_{t})+\frac{\alpha}{r}M_{0}\right)$
+$=\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\left(1-\frac{\alpha}{r}\right)\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{w}}_{t})$
 $\leq\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t}).$
@@
 $\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}+\left(\frac{6\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T.$
-To prove the result in the Lipschitz case, we note that, according to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$.
+To prove the result in the Lipschitz case, we note that, according to Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct, we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$.
 By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
@@
 On the other hand, we have
-$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
-$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
-$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
-$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\alpha}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
-$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\alpha M_{1}D}{r}\right)$ (Lipschitz)
+$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{w}}}_{t})$
+$\leq\delta M_{1}T+\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{w}}}_{t})$ (Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct)
+$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{w}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
+$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{w}}_{t}+\frac{\alpha}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
+$\leq\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{w}}_{t})+\frac{2\alpha M_{1}D}{r}\right)$ (Lipschitz)
 $=\left(1+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
@@
 Proof of Theorem 10.
-The proof of the bounds for this case is similar to the previous case.
-As before, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
-It follows from the description of Meta-algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+[…]
 Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$.
@@
 $={\mathbb{E}}\left[\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})\right)+\left(\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right].$ (8)
-To obtain the same bound as before, we note that in the convex case, instead of Inequality 8, we have
+To obtain the same bound as before, we note that in the convex case, instead of Inequality 6, we have
 $|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\frac{2M_{0}\delta}{\alpha}<\frac{4M_{0}\delta}{\alpha},$ (9)
-and, in the Lipschitz case, instead of Inequality 9, we have
+and, in the Lipschitz case, instead of Inequality 7, we have
 $|f_{t}({\mathbf{x}}_{t})-\hat{f_{t}}({\mathbf{x}}_{t})|\leq\delta M_{1}<2\delta M_{1}.$
 The rest of the proof follows verbatim. ∎
-G Proof of Theorem 8
-This proof is similar to the proofs of Theorems 6 and 7.
+Appendix G Proof of Theorem 14
+This proof is similar to the proofs of Theorems 10 and 11.
 Proof.
@@
 where $\mathbf{v}$} is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$.
-Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Lemma 5 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.
+Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Lemma 5 in pedramfar23_unified_approac_maxim_contin_dr_funct, $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.
 Moreover, we have
-$\|\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}}\|\leq\frac{k}{2\delta}M_{1}\|({\mathbf{x}}+\delta{\mathbf{v}})-({\mathbf{x}}-\delta{\mathbf{v}})\|\leq kM_{1}.$
-Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},kM_{1})$.
-Using Equation 12 and the definition of the Meta-algorithm 4, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
-It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
+$\left\|\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}}\right\|\leq\frac{k}{2\delta}M_{1}\|({\mathbf{x}}+\delta{\mathbf{v}})-({\mathbf{x}}-\delta{\mathbf{v}})\|\leq kM_{1}.$
+[…]
 Let ${\mathbf{u}}\in\operatorname{argmin}_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$ and $\hat{{\mathbf{u}}}\in\operatorname{argmin}_{{\mathbf{u}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{u}}_{t})$.
@@
 $={\mathbb{E}}\left[\left(\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\sum_{t=a}^{b}\hat{f_{t}}({\mathbf{x}}_{t})\right)+\left(\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})-\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right)\right].$ (11)
-According to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. On the other hand, we have
-$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
-$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
-$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\delta}{r}\right){\mathbf{u}}_{t}+\frac{\delta}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
-$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\delta}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
-$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\delta M_{1}D}{r}\right)$ (Lipschitz)
+According to Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct, we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. On the other hand, we have
+$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{w}}}_{t})$
+$\leq\delta M_{1}T+\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{w}}}_{t})$ (Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct)
+$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\delta}{r}\right){\mathbf{w}}_{t}+\frac{\delta}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
+$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{w}}_{t}+\frac{\delta}{r}({\mathbf{c}}-{\mathbf{w}}_{t})\right)$
+$\leq\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{w}}_{t})+\frac{2\delta M_{1}D}{r}\right)$ (Lipschitz)
 $=\left(1+\frac{2D}{r}\right)\delta M_{1}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$